### v1

# ReAgent: Reversible Multi-Agent Reasoning for Knowledge-Enhanced Multi-Hop QA

Abstract
Recent advances in large language models (LLMs) have significantly improved multi-hop question answering (QA) through direct Chain-of-Thought (CoT) reasoning. However, the irreversible nature of CoT leads to error accumulation, making it challenging to correct mistakes in multi-hop reasoning. This paper introduces ReAgent: a Reversible multi-Agent collaborative framework augmented with explicit backtracking mechanisms, enabling reversible multi-hop reasoning. By incorporating text-based retrieval, information aggregation and validation, our system can detect and correct errors mid-reasoning, leading to more robust and interpretable QA outcomes. The framework and experiments serve as a foundation for future work on error-tolerant QA systems. Empirical evaluations across three benchmarks indicate ReAgent’s efficacy, yielding average about 6% improvements against baseline models.
ReAgent: Reversible Multi-Agent Reasoning for Knowledge-Enhanced Multi-Hop QA
Zhao Xinjie1, Fan Gao1, Rui Yang2, Yingjian Chen1, Yuyang Wang3, Ying Zhu1, Jiacheng Tang1, Irene Li1 1 The University of Tokyo, 2 Duke-NUS Medical School, 3 The Hong Kong University of Science and Technology (Guangzhou)
1 Introduction
Multi-hop question answering (QA) represents a crucial challenge in natural language processing (NLP), requiring systems to integrate and reconcile evidence from multiple sources or hops of reasoning before reaching a single final answer (Yang et al., 2018; Welbl et al., 2018; Ho et al., 2020). Existing benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020) have shed light on the intrinsic difficulty of multi-hop inference: not only must relevant information be located across dispersed documents or knowledge graphs, but the process itself must remain robust to partial mistakes. If an early inference is incorrect, subsequent steps can amplify that error and produce an unreliable or contradictory final answer (Inoue et al., 2020; Bo et al., 2024; Yang et al., 2024c).
Recent large language models (LLMs) have showcased strong performance on multi-hop QA, often employing either explicit Chain-of-Thought (CoT) prompting (Wang et al., 2023) or retrieval-augmented generation (Das et al., 2018; Long et al., 2019; Gao et al., 2023; Liu et al., 2023; Yang et al., 2024b, 2025). While these strategies can help track intermediate steps, they typically rely on a predominantly forward-directed workflow. Once a partial conclusion is formed, the system rarely revisits or corrects it unless externally prompted (Puerto et al., 2023; Huang et al., 2024). This forward-only structure poses limitations, especially when later evidence contradicts an earlier inference or when multiple lines of reasoning yield divergent findings. A single incorrect hop can invalidate the final answer, and without explicit backtracking, the system is unable to retract and repair flawed assumptions (Doyle, 1979; Bo et al., 2024).
Figure 1: We compare different frameworks for QA: Linear (CoT), Collaborative (MAS), and Reversible (ReAgent) Reasoning in Multi-hop QA.
To address this gap, growing attention has been directed to multi-agent collaboration for QA, where specialized agents focus on sub-tasks—such as retrieval, verification, and resolution of conflicting information—and then synchronize their outcomes (Zhao et al., 2024; Parhizkar et al., 2020; Ke et al., 2024). By delegating different roles (e.g., question decomposition, result verification) to different agents, the system becomes more modular and potentially more explainable. Recent multi-agent systems include COPPER (Bo et al., 2024) and LongAgent (Zhao et al., 2024), both of which decompose complex tasks among multiple LLM-based agents that coordinate via message passing. They not only reduce overall context per agent but can also cross-check for consistency, flagging suspicious partial solutions. However, even these advanced frameworks often lack fully reversible reasoning pipelines, wherein each agent can effectively backtrack to an earlier valid state once a conflict is detected, re-explore alternatives, and propagate the corrections forward (He et al., 2021).
In this paper, we introduce ReAgent, a reversible multi-agent collaborative reasoning framework with explicit backtracking, designed for multi-hop question answering. This approach builds on the premise that in real-world complex QA scenarios—especially when reasoning across a broad knowledge graph or multiple textual sources—one-size-fits-all steps fail to capture the frequent need for partial re-evaluation (Yang et al., 2018; Ho et al., 2020). Conflict detection and backtracking mechanisms act as a flexible error-correction loop, in which a mismatch in evidence can trigger a re-check of previously accepted facts, ensuring that misinterpretations do not survive to the final answer. This perspective aligns with “reflective” multi-agent designs (Bo et al., 2024; Huang et al., 2024), but pushes further by defining explicit multi-layer protocols—local versus global backtracking—and employing a dynamic aggregator that leverages trust signals from each agent’s past reliability (Puerto et al., 2023; Parhizkar et al., 2020).
We empirically show that this reversible approach improves final-answer accuracy and interpretability: backtracking traces clarify how the system isolates and corrects mistakes. Our work links robust non-monotonic reasoning with data-driven QA, offering a transparent, error-tolerant multi-hop inference pipeline. In summary, our main contributions are: (1) a novel multi-agent QA framework that supports backtracking at both local and global levels; (2) hybrid retrieval from knowledge graphs and text for flexible evidence acquisition; and (3) systematic experiments on three benchmark datasets demonstrating more robust and interpretable multi-hop QA than forward-only baselines. Specifically, our framework achieves an average improvement of approximately 6% over baseline models.
Figure 2: Execution, Supervision, and Interaction Layers: Three-Fold Role Partitioning in ReAgent’s Multi-Hop QA Backtracking Mechanisms.
2 Preliminary
[…]
Challenges in Forward-Only Reasoning A key source of error in multi-hop QA lies in the lack of revision capacity. Once a system commits to a specific inference early on (for instance, deducing a partial conclusion from $e_{1}$ and $e_{2}$), there is no built-in mechanism to retract or update that inference if additional evidence suggests it is incorrect. As the system progresses to hop $j{+}1$ or further, it carries potential inaccuracies along, making the final answer unreliable or inconsistent. In many cases, the model would have benefited from re-examining initial assumptions, especially when encountering contradictory or low-confidence signals. These observations motivate an interactive and flexible approach, wherein the QA process is allowed to backtrack and replace problematic intermediate results.
Non-Monotonic Backtracking Our notion of backtracking is a non-monotonic extension of the typical multi-hop QA process. Rather than treating the knowledge state as strictly accumulative, we allow for partial or global reversion. Specifically:
Local Correction: An agent can retract one of its own inferences if it detects internal conflicts or if it receives contradictory evidence from other agents.
Global Rollback: When inconsistencies span multiple agents, a supervisory mechanism coordinates a synchronized rollback. This can mean discarding a sequence of accepted statements that are ultimately untenable, returning the system to a previously consistent state.
This extended capability is guided by two key considerations. First, not all new evidence is guaranteed to be more credible than old; hence, the system must weigh or prioritize contradictory claims. Second, if an inference $I$ made in hop $r$ is discredited in hop $r{+}2$, it may be more efficient to rollback to hop $r$ directly, rather than forcing the pipeline to proceed with a known contradiction. Such rollback actions effectively prevent one faulty hop from invalidating the entire chain of reasoning.
Figure 3: Layered Multi-Agent Interaction and Backtracking Flow with Time Step.
3 ReAgent: Reversible Multi-Agent Reasoning Architecture
3.1 Overview of the ReAgent Architecture
Figure 2 presents the overall design, organized into three layers. Let $\mathcal{A}\;=\;\{A_{1},A_{2},\dots,A_{n}\}$ be the set of agents. At time step $t$, each agent $A_{i}$ maintains a local knowledge set $\Phi_{i}^{t}$ that may include proposed assertions, retrieved evidence, or partial inferences. In addition, the system can track a global assertion set $\Phi^{t}\;=\;\bigcup_{i=1}^{n}\Phi_{i}^{t},$ representing all statements under consideration at time $t$. The three-layer structure is as follows:
  1. 1.
    Execution Layer (Section 3.2): Comprises the essential agents for domain tasks, such as decomposing questions, retrieving evidence, verifying local consistency, and assembling intermediate answers.
  2. 2.
    Supervisory Layer (Section 3.3): Includes supervisory agents that orchestrate conflict resolution and manage global backtracking, ensuring that contradictory or invalid assertions are addressed across all relevant agents.
  3. 3.
    Interaction Layer (Section 3.4): Maintains the concurrency model and communication protocols. It defines how agents send and receive messages (e.g., assert, reject, challenge) and how the system merges these parallel events into consistent updates of $\Phi^{t}$.
By separating concerns across these layers, it becomes simpler to reason about local agent updates while still allowing global interventions. The design explicitly supports non-monotonic updates: newly introduced statements can be revoked if they lead to logical conflicts or are superseded by contradictory evidence. The specific prompts of each kind of agent is provided in Appendix A.
3.2 Execution Layer Agents
The Execution Layer hosts four types of agents that address fundamental QA subtasks, each maintaining its local knowledge $\Phi_{i}$. When necessary, an agent may perform local backtracking to remove or adjust recent assertions.
Question-Decomposer Agent ($A_{Q}$).
Breaks a complex input question $Q_{0}$ into sub-questions $\{q_{1},q_{2},\dots\}$. If $Q_{0}(x)$ denotes the main query, $A_{Q}$ produces a decomposition mapping $\{q_{j}\}$ for subsequent retrieval and verification: $A_{Q}:\;Q_{0}(x)\;\longmapsto\;\{\,q_{1},\dots,q_{m}\}.$ This decomposition is broadcast to other base-layer agents.
Retriever Agent ($A_{R}$).
Upon receiving a sub-question $q_{i}$, $A_{R}$ returns relevant evidence $E$ (e.g., text passages or knowledge graph triples), possibly with confidence scores $\alpha(E)$. This agent may maintain a local retrieval cache and re-invoke searches if backtracking invalidates prior evidence.
Verifier Agent ($A_{V}$).
Performs local consistency checks. Given new evidence $E$, it verifies coherence with $\Phi_{V}^{t}$ (the Verifier’s current knowledge). If conflicts arise, $A_{V}$ triggers local backtracking to revert to a consistent state or signals higher-level agents for broader resolution. Concretely, $A_{V}$ may invoke
$\mathrm{BacktrackLocal}\bigl{(}A_{V},r\bigr{)}$
if it suspects that assertions introduced after node $r$ are inconsistent.
Answer-Assembler Agent ($A_{A}$).
Gathers partial answers (and verified evidence) to synthesize a final solution. If the set of local inferences from $A_{Q}$, $A_{R}$, and $A_{V}$ is $\Phi_{\text{base}}$, then $A_{A}$ combines them to generate the final answer $A_{\text{final}}$. Any detected contradiction triggers further checks or possible escalation to the supervisory layer.
3.3 Supervisory Layer Agents
The supervisory layer manages conflicts that cannot be resolved by a single agent. It also oversees system-wide strategies, especially when contradictory goals or inconsistent states appear across multiple agents.
Supervisor Agent ($A_{S}$).
Coordinates multi-agent conflicts spanning multiple knowledge sets or critical shared assumptions. If a conflict persists after local backtracking, $A_{S}$ determines whether partial or holistic rollback is required. When a system-wide contradiction is found, one may identify a minimal conflict set $\Psi$ such that $\mathrm{SAT}(\Psi)=\textsf{false}$ but $\mathrm{SAT}(\Psi^{\prime})=\textsf{true}$ for any proper subset $\Psi^{\prime}\subset\Psi$. The supervisor then triggers a backtracking operation that eliminates or modifies $\Psi$.
Controller Agent ($A_{C}$).
Regulates high-level strategies, monitoring game-theoretic signals or meta-rules. If certain rules appear suboptimal or unsafe, $A_{C}$ can override them or enforce mode switches. For instance, it may issue a $\textsf{challenge}(\varphi)$ to examine a crucial assertion $\varphi$ from multiple agents’ perspectives, or override local decisions if they jeopardize overall consistency.
3.4 Interaction Layer
Communication Protocols.
We adopt a message-driven framework, where atomic events $\mathtt{msg}(\varphi)$ are labeled with types such as assert, inform, reject, or challenge. These updates can occur simultaneously, requiring a composite event model to merge parallel messages. A shared concurrency mechanism processes simultaneous actions, ensuring that if conflicts arise from concurrent assertions, they are escalated to the supervisory layer.
Temporal Operators.
The system can utilize operators like $\square\varphi$ to denote that $\varphi$ must hold in every future state, and $\lozenge\varphi$ to indicate that a proposition $\varphi$ might become true at some future point. These temporal constraints assist in specifying persistent axioms or potential triggers for backtracking conditions.
Algorithm 1 Multi-Agent Reversible Reasoning
1:$Q_{0}$: Main question
$A_{Q}.\mathrm{decompose}(Q_{0})\to\{q_{1},q_{2},\dots\}$
3:Broadcast $\mathrm{assert}(q_{i})$ to $A_{R},A_{V},A_{A}$
4:for each $q_{i}$ in $\{q_{1},q_{2},\dots\}$ concurrently do
5:     $E\leftarrow A_{R}.\mathrm{retrieve}(q_{i})$
6:     $A_{V}.\mathrm{verify}(E)$
7:     if $A_{V}$ detects conflict locally then
8:         $A_{V}.\mathrm{BacktrackLocal}(r_{V})$
9:         if conflict persists then
10:              raise Conflict to $A_{S}$
11:         end if
12:     end if
13:     $A_{A}.\mathrm{storePartialAnswer}(q_{i},E)$
14:end for
15:if Global conflict signaled then
16:     $A_{S}.\mathrm{HolisticUpdate}(\Pi_{j})$
17:     if $A_{C}$ intervention needed then
18:         $A_{C}.\mathrm{challenge}(\varphi)$ or override
19:     end if
20:end if
21:Final $\leftarrow A_{A}.\mathrm{assembleAnswer}(\{q_{1},q_{2},\dots\})$
22:return Final Answer
3.5 Conflict Management and Backtracking
This part formalizes how the system detects contradictions and reverts to consistent states. At each time $t$, the set of all accepted assertions is $\Phi^{t}\;=\;\bigcup_{i=1}^{n}\Phi_{i}^{t}.$ A conflict occurs if $\Phi^{t}$ entails both $\phi$ and $\neg\phi$ for some proposition $\phi$, meaning $\mathrm{SAT}(\Phi^{t})$ fails. To resolve the issue, the system follows a two-level approach:
Local Backtracking.
Each agent $A_{i}$ maintains a backtracking graph $\mathrm{LBG}_{i}$, which logs states $\Phi_{i}^{r}$ at selected checkpoints $r$. If $\mathrm{Conflict}_{i}(\Phi_{i}^{t})$ indicates an internal contradiction, $A_{i}$ tries to revert to a prior consistent node $r<t$:
$\mathrm{BacktrackLocal}(A_{i},\,r):\quad\Phi_{i}^{t}\;\longrightarrow\;\Phi_{i}^{r}.$
After local rollback, the agent re-evaluates newly arriving evidence.
Global Conflict and Holistic Updates.
When contradictions span multiple agents, the Supervisor Agent identifies a minimum conflict set of assertions that must be revised. A global backtracking operation
$\mathrm{BacktrackGlobal}\bigl{(}\sigma^{t},\,r\bigr{)}$
reverts all agents from time $t$ to $r$. The system discards any statements introduced between $r+1$ and $t$, restoring consistency. If the conflict cannot be removed even after a global rollback, the Controller might enforce strategic overrides or disclaim an answer.
Algorithm 1 summarizes the flow of this multi-agent reversible reasoning design. The process starts with question decomposition and sub-question retrieval, followed by verification. Local backtracking is invoked as needed to address internal inconsistencies. If the conflict persists, it is escalated to the supervisory layer, where partial or holistic backtracking may be applied. After conflicts are resolved, the final step integrates all consistent sub-answers.
Model HotpotQA 2Wiki Musique Average
EM F1 EM F1 EM F1 EM F1
Zero-shot
Qwen-2.5-14b-Instruct 0.205 0.316 0.258 0.308 0.042 0.131 0.168 0.251
Qwen-2.5-32b-Instruct 0.191 0.304 0.190 0.294 0.041 0.133 0.140 0.243
Qwen-2.5-72b-Instruct 0.238 0.355 0.286 0.338 0.049 0.141 0.191 0.278
Llama-3.3-70b-Instruct 0.280 0.415 0.290 0.349 0.063 0.155 0.211 0.306
DeepSeek-v3 0.293 0.410 0.258 0.332 0.042 0.145 0.197 0.295
Gemini-1.5-flash 0.236 0.348 0.267 0.311 0.055 0.148 0.186 0.269
Gemini-2.0-flash 0.317 0.434 0.335 0.391 0.075 0.170 0.242 0.331
GPT-4o 0.359 0.497 0.346 0.417 0.100 0.217 0.268 0.377
Knowledge Augmented
Qwen-2.5-14b-Instruct 0.594 0.768 0.623 0.725 0.305 0.453 0.507 0.648
Qwen-2.5-32b-Instruct 0.576 0.756 0.537 0.694 0.324 0.460 0.479 0.636
Qwen-2.5-70b-Instruct 0.604 0.772 0.664 0.753 0.326 0.475 0.531 0.667
Llama-3.3-70b-Instruct 0.624 0.790 0.698 0.783 0.332 0.470 0.551 0.681
DeepSeek-v3 0.620 0.792 0.673 0.769 0.368 0.505 0.553 0.688
Gemini-1.5-flash 0.622 0.784 0.644 0.732 0.314 0.435 0.526 0.680
Gemini-2.0-flash 0.623 0.790 0.638 0.744 0.306 0.441 0.522 0.685
GPT-4o 0.612 0.794 0.677 0.780 0.356 0.506 0.548 0.693
ReAgent (Ours) 0.630 0.795 0.711 0.793 0.371 0.515 0.571 0.701
Table 1: Performance of different models across three multi-hop QA datasets. In each column, the highest performance is highlighted in bold; and within each method group, the top performer is underlined.
4 Experiments
4.1 Experimental Setup
We assessed ReAgent through three widely used knowledge-intensive multi-hop QA datasets: HotpotQA (Yang et al., 2018), 2WikiMultiHotQA (Ho et al., 2020) and Musique (Trivedi et al., 2022). We show how ReAgent produces better results compared to baseline models and effectively solves problems requiring continuously corrected reasoning.
Datasets
HotpotQA is a large-scaled, open-domain dataset encouraging multi-hop reasoning. The dataset provides full Wikipedia passages for information retrieval and integration. 2WikiMultiHopQA selects distinct Wikipedia domains for each question, focusing on evaluating the model’s multi-hop inference over different resources rather than one document. Musique is a challenging dataset requiring model’s ability to reason over multiple dispread sentences.
Following previous work Trivedi et al. (2023); Press et al. (2023); Gutiérrez et al. (2024), we utilize 1000 random samples from each validation set and corresponding texts as the knowledge base.
Baselines.
We include both proprietary and open-sourced LLMs with multiple sizes for evaluation: Qwen-2.5 Yang et al. (2024a) of varying parameter scales, the newly Llama-3.3 Meta (2024) and DeepSeek-v3 Liu et al. (2024), Gemini Team et al. (2023) variants (Gemini-1.5-flash-2024-09, Gemini-2.0-flash-2025-02 as well as GPT-4o-2024-08-06 Hurst et al. (2024).
Settings.
We hold two distinct settings for the performance evaluation: Zero-shot with CoT only provides the question, expecting to validate LLM’s performance with internal knowledge; Knowledge-Augmented with CoT Liang et al. (2024); Gutiérrez et al. (2024) setting introduces related documents for reasoning, evaluating model’s capability of information integration especially processing multiple mixed evidence.
Implementation.
For super large-size LLMs (DeepSeek-v3, Gemini families, GPT-4o), we access through API, for medium-size LLMs, we access from their official repositories. We set the temperature as $0.2$ to get deterministic answers as well as somehow maintain flexibility. We utilize the default settings of other parameters. ReAgent utilizes GPT-4o as the backbone for execution. Specifically, we set the temperature for decomposition agent as $0.8$, the remaining as $0.6$. All prompts are available in Appendix A.
Metrics.
We measure Exact Match (EM) and F1 scores for the QA evaluation.
4.2 Results and Analysis
Table 1 illustrates the comprehensive performance. ReAgent outperforms all the baseline models in both metrics across all datasets. Specifically, our framework achieves an average EM of 0.571 and an average F1 of 0.701, outperforming GPT-4o by 2.3% in EM and 0.8% in F1.
External Knowledge Matters.
Definitely, we observe that zero-shot performance is substantially lower, reflecting the difficulty of multi-hop reasoning without relevant background. With intensive knowledge, all models outperform their zero-shot versions in different levels, presenting their internal reasoning abilities.
Scaling-law not always works.
We carefully select baseline models across small(14b), medium(70b), large(>100b) size for comparisons. Interestingly, the results show that smaller models sometimes outperform larger models. For example, Qwen-2.5-14b model generally defeats its 32b counterparts in both zero-shot and knowledge augmented settings. Similarly, Llama-3.3 approach or even surpass the super large-scale models (e.g. Gemini-1.5-flash and DeepSeek-v3 in zero-shot, Gemini-2.0-flash and GPT-4o in knowledge-intensive setting). This suggests that multi-hop QA can be aided by carefully tuned architectures, especially if they incorporate chain-of-thought or knowledge-aware strategies.
Overall, we emphasized the value of reversible reasoning mechanisms to mitigate error propagation. The data in Table 1 align with this assumption: even with strong models, single-pass reasoning can fail unless the correct context is identified or re-checked. Our method’s ability to backtrack helps resolve contradictions, leading to more stable performance in multi-hop scenarios.
Figure 4: Ablation Study Results Comparison.
4.3 Ablation Study
To gauge the importance of the backtracking mechanism, we conduct an ablation study by disabling it under the same settings. Specifically, the system then proceeds in a strictly forward fashion, without the ability to revert to earlier intermediate states. Fig.4 compares performance between different backbones, with and without backtracking, on three datasets. As shown, the performance drop highlights the importance of backtracking in mitigating error propagation, where an early misstep without backtracking cannot be corrected, resulting in deteriorated performance. Notably, the DeepSeek-v3 backbone exhibits general improvements with backtracking, demonstrating its robustness across various settings.
5 Case Study
We illustrate our multi-agent backtracking (ReAgent) through a single-culprit puzzle with four suspects {A, B, C, D}. Each suspect gives statements about who might be guilty or lying. The puzzle’s only correct solution is that B is the culprit, but identifying this requires partially retracting initial assumptions along the way. Detailed puzzle statement can be found in Appendix B.
Why a Rollback Mechanism is Needed.
Naive single-pass (or single-thread) models often fixate prematurely on one suspect, disregard contradictory evidence, and produce unsound conclusions. By contrast, our multi-agent approach enumerates possible culprits in parallel, detects conflicts, and rolls back to revise incorrect assumptions.
Below is a simplified agentic workflow capturing the time steps $T_{1}$ through $T_{7}$ during which each agent processes statements, detects a conflict, and reverts.
Algorithm 2 ReAgent Multi-Agent Puzzle Solving
1:Input: Puzzle statements from A, B, C, D.
2:Goal: Identify unique culprit (one of {A,B,C,D}) with minimal contradictions.
3:Initialization
[…]
6:Round 1: Checking A as culprit
7:$T_{1}:$ $\text{Checker}_{A}$:  Evaluate puzzle statements assuming $A$ is culprit.
8:$T_{2}:$ $\text{Checker}_{A}$ detects contradiction:  ($A$’s claims vs. $C$’s claims cannot both hold).
9:$T_{3}:$ Conflict Detector signals rollback to discard $A$-culprit assumption.
10:Round 2: Checking B as culprit
11:$T_{4}:$ $\text{Checker}_{B}$ collects statements $\{A,B,C,D\}$ under $B$=culprit.
12:$T_{5}:$ Finds no fatal conflicts (B’s statements can be partly false, others partly true).
13:$T_{6}:$ Conflict Detector sees consistency, no rollback needed.
14:Rounds 3 and 4: Checking C or D as culprit
15:$T_{7}:$ Similar to $H_{A}$, each leads to irreconcilable contradictions.
16:          $\Rightarrow$ Supervisor triggers rollback again; discards these.
17:Conclusion:  Only $H_{B}$ remains consistent throughout.  Answer: $B$ is culprit.
Discussion.
This staged procedure allows each culprit hypothesis to be explored (lines 7–13 of Algorithm 2). Any contradictions are escalated at $T_{2}$ or $T_{7}$ to the Conflict Detector, which triggers a rollback to the prior consistent node. Ultimately, only the assignment “$B$ is culprit” withstands these iterative checks. Without rollback, a linear solver often locks in a wrong suspect early and fails to correct itself. For extended discussions of single-thread LLM failures, and the complete multi-agent transcripts, see Appendix B and  C.
6 Related Work
6.1 Prompting and Iterative Reasoning
Large Language Models (LLMs) can solve complex tasks by leveraging prompting strategies that encourage step-by-step reasoning, often referred to as chain-of-thought (CoT) prompting Wei et al. (2022), which has demonstrated improved performance in arithmetic, commonsense, and symbolic reasoning tasks, particularly in sufficiently large models. Sampling multiple reasoning paths and selecting the most consistent answer, known as self-consistency, further boosts accuracy by treating final answers as a “majority vote” across different solutions Wang et al. (2023). To reduce the need for hand-tuned prompts, automated prompt engineering techniques have been explored. One approach, Automatic Prompt Engineer (APE), treats prompt generation as a search problem: one LLM proposes candidate instructions while another evaluates them, allowing the system to refine prompts over multiple tasks Zhou et al. (2023). LLMs can thus act as both reasoning engines and prompt designers. Beyond one-shot responses, iterative refinement frameworks empower LLMs to correct and enhance their outputs. For instance, the ReAct framework interleaves reasoning traces with actions (e.g., external API calls), enabling LLMs to incorporate feedback and reduce factual errors Yao et al. (2023). Other methods like Reflexion Shinn et al. (2023) and Self-Refine Madaan et al. (2023) introduce self-feedback loops, where a model critiques its own response and revises it in subsequent attempts, leading to more reliable and transparent solutions. Prompting methods have also expanded into multimodal settings, where models process both text and visual inputs. In Multimodal-CoT, the system first generates a textual chain-of-thought integrating visual information, then produces an answer from that combined rationale Zhang et al. (2023).
6.2 Multi-Agent Collaboration, Debate, and Scalability
Recent work explores multi-agent approaches, where multiple LLMs (or instances of the same model) collaborate or compete toward a shared goal. For instance, the CAMEL framework assigns different roles (e.g., “user” and “assistant”) to LLM agents that communicate in a dialogue-based fashion, breaking down tasks more systematically Li et al. (2023). Role specialization allows agents to focus on different aspects of the problem, and voting or consensus mechanisms—akin to self-consistency—can yield more robust answers. Debate-oriented methods push multi-agent interactions further by having agents argue for different viewpoints. After hearing opposing arguments, a human or model judge selects the most convincing solution Khan et al. (2024). Similarly, in an interconsistency framework, multiple agents provide arguments and counterarguments over multiple rounds, eventually reaching a consensus Xiong et al. (2023). Such adversarial or critical exchanges can improve truthfulness and correctness, though an overly dominant model can bias outcomes unless moderated by a sufficiently capable judge Xiong et al. (2023). These multi-agent and multimodal systems introduce scalability and coordination challenges. As more agents or steps are added, communication overhead can grow significantly, making real-time collaboration difficult. Research highlights the need for efficient structures—e.g., hierarchical controllers or learned protocols—to balance the benefits of specialized agents with the system’s complexity Li et al. (2023); Xiong et al. (2023).
7 Conclusion
In this paper, we introduced a multi-agent QA framework, ReAgent, that incorporates backtracking mechanisms to mitigate error propagation in multi-hop reasoning. Our hierarchical approach addresses the long-standing challenge that a single misstep during inference often invalidates the entire reasoning chain. By allowing partial or global rollback, each agent can detect and correct conflicting evidence, leading to more stable and interpretable solutions. Experiments on three multi-hop QA benchmarks—HotpotQA, 2WikiMultiHopQA, and Musique—demonstrate that explicit backtracking and conflict resolution improve performance beyond forward-only baselines, confirming the importance of error revision for complex question answering. We believe that the reversible, modular design can be extended to other knowledge-intensive applications, laying the groundwork for more trustworthy collaborative AI agents.
Limitations
Despite its advantages, our framework inherits several constraints:
  • Complex Coordination: Multi-agent interaction and backtracking mechanisms introduce additional computational overhead and higher implementation complexity.
  • Dependence on Retrieval Quality: As with many QA systems, our accuracy hinges on retrieving relevant and sufficiently detailed evidence. Misleading or scarce information can impair performance, even with backtracking.
  • Scalability to Larger Contexts: While our approach manages conflicts through localized rollback, handling very large or highly diverse corpora may require more sophisticated retrieval and aggregation strategies to maintain efficiency.
  • Model Assumptions: We rely on the underlying language models’ capacity for reliable verification. If the base model’s reasoning is systematically biased, backtracking alone may not fully correct errors.
Ethic Statement
Our method focuses on improving the interpretability and robustness of multi-hop question answering rather than amplifying any content that could be personally or socially harmful. We do not incorporate user-sensitive or private data, and the knowledge sources employed are publicly accessible. However, as with any automated system, inaccurate or biased inputs may lead to flawed inferences if not properly filtered or monitored. We encourage careful data curation and ongoing human oversight to mitigate unintended consequences. Furthermore, we recommend that users verify critical or sensitive outputs, since backtracking techniques reduce—but cannot entirely eliminate—residual errors inherent in large language models.
Appendix
Appendix A Multi-Agent Prompt Templates
A.1 Question-Decomposer Agent
Question-Decomposer Agent Prompt Role Description: You are the Question-Decomposer Agent, specializing in breaking down the user’s complex query into smaller, manageable sub-questions or sub-tasks. This decomposition is crucial for multi-hop question answering and will be consumed by downstream agents (Retriever, Verifier, etc.) in the pipeline. Your Goals: 1. Parse the original query into logically independent or sequential sub-questions. 2. Preserve all necessary context so that other agents can retrieve relevant evidence and validate partial answers. 3. Output your decomposition in a structured JSON format. Example Usage: Original Query: “Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?” Decomposition: $q_{1}$: Identify which U.S. state hosted the 1984 Summer Olympics. $q_{2}$: Find the capital city and the largest city of that state. $q_{3}$: Compare population sizes of the capital and largest city. $q_{4}$: Return the state if the capital is indeed smaller. Output Format (JSON Only): { "sub_questions": [ "Sub-question 1", "Sub-question 2", ... ], "decomposition_reasoning": "A short textual explanation of your decomposition process" } Instruction: Please ensure your final output is a valid JSON object matching the above schema.
A.2 Retriever Agent
Retriever Agent Prompt Role Description: You are the Retriever Agent, responsible for fetching relevant evidence from external sources (a text corpus, a knowledge graph, or both) based on sub-questions provided by the Question-Decomposer Agent. This includes documents, passages, knowledge graph triples, and any other data needed for multi-hop QA. Your Goals: 1. Given a sub-question, retrieve the most relevant facts or passages. 2. Include confidence scores or other metadata if available. 3. Return your findings in a standardized JSON structure so that the Verifier and Answer-Assembler Agents can process them. Example Usage: Input Sub-question: “Which U.S. state hosted the 1984 Summer Olympics?” Retrieved Evidence (text-based): { "document": "History of the Olympics", "passage": "The 1984 Summer Olympics were held primarily in Los Angeles, California." } Output Format (JSON Only): { "retrieved_evidence": [ { "source": "e.g., ’Wikipedia excerpt’ or ’KG triple ID’", "content": "string or structured data relevant to the sub-question", "confidence": 0.0-1.0 (optional) }, ... ], "retrieval_reasoning": "Short justification for why this evidence is relevant" } Instruction: Output only valid JSON, strictly matching the above schema.
A.3 Verifier Agent
Verifier Agent Prompt Role Description: You are the Verifier Agent, focusing on assessing consistency and correctness of the newly retrieved evidence or intermediate inferences. You detect contradictions or conflicts either within the new data or against the previously verified knowledge. If necessary, you trigger local backtracking to remove or adjust statements causing inconsistency. Your Goals: 1. Validate whether new information is consistent with existing verified knowledge. 2. Identify contradictions and either correct them or escalate them to a higher-level supervisor if unresolved. 3. Produce a final set of verified facts or a signal indicating a conflict. Example Usage: Incoming Evidence: "Sacramento population: 508,000" and "Sacramento population: 1,500,000" Detected Inconsistency: "Sacramento cannot have two drastically different population values." Local Backtracking Action: "Discard the erroneous or lower-confidence figure (1,500,000)." Output Format (JSON Only): { "verified_facts": [ "Fact 1", "Fact 2", ... ], "conflicts_detected": [ "Conflict 1 description, if any", ... ], "local_backtracking_action": "Description of any backtracking performed, or ’none’" } Instruction: Return only valid JSON with the fields above. Provide a concise summary if backtracking occurs.
A.4 Answer-Assembler Agent
Answer-Assembler Agent Prompt Role Description: You are the Answer-Assembler Agent. You gather verified facts from the Verifier Agent and partial answers from the Decomposer and Retriever Agents. You then synthesize a coherent, contextually relevant response, producing a final or intermediate answer for the user. If you detect a major conflict among partial answers, you escalate to the Supervisor Agent. Your Goals: 1. Aggregate partial answers logically. 2. Compose a natural-language (or structured) final answer to the user’s multi-hop query. 3. Escalate unresolvable contradictions to the Supervisor Agent if needed. Example Usage: Partial Answers and Verified Facts: {"hosted_1984_olympics": "California"}
{"capital": "Sacramento"}, {"largest_city": "Los Angeles"}
{"pop_sacramento": 508000}, {"pop_los_angeles": 3900000}
Composed Final Answer: "The state is California, since its capital city (Sacramento) has a population smaller than that of Los Angeles."
Output Format (JSON Only): { "final_answer": "A concise or structured answer to the main query", "partial_answer_synthesis": [ "Short bullet points on how partial answers were combined" ], "escalation_signal": "Set to ’none’ if no escalation is needed, otherwise a short reason" } Instruction: Return only valid JSON. If you detect a major conflict you cannot resolve, set "escalation_signal" to a reason for Supervisor Agent intervention.
A.5 Supervisor Agent
Supervisor Agent Prompt Role Description: You are the Supervisor Agent, responsible for orchestrating global conflict resolution and global backtracking if needed. When partial answers or verified facts across multiple agents yield irreconcilable contradictions, you identify a minimal conflict set and roll back the entire system’s state to a previously consistent checkpoint if local fixes fail. Your Goals: 1. Collect escalation signals from the Answer-Assembler or Verifier Agents. 2. If local backtracking does not resolve the conflict, execute system-wide or multi-agent rollback. 3. Provide a summary of changes, indicating which statements or partial answers are discarded or revised. Example Usage: Received Escalation: “Capital(California, Sacramento) conflicts with Capital(California, Los Angeles).” Global Backtracking Action: “Rollback to a state before the second capital claim was introduced. Discard that erroneous claim from the knowledge base.” Output Format (JSON Only): { "conflict_summary": [ "Brief descriptions of contradictory sets" ], "global_backtracking_decision": "Description of how far to roll back or ’none’", "updated_consensus_state": [ "Any statements or facts that remain accepted after rollback" ], "reasoning_notes": "Explanation of the chosen resolution strategy" } Instruction: Output valid JSON. If no global conflict is found, indicate "none" for "global_backtracking_decision".
A.6 Controller Agent
Controller Agent Prompt Role Description: You are the Controller Agent, providing high-level strategic oversight. You may override local decisions, impose extra checks, or challenge specific assumptions if repeated conflicts persist. You also maintain meta-information such as agent reliability scores or fallback strategies. Your Goals: 1. Intervene in situations where standard local or global backtracking repeatedly fails. 2. Challenge or confirm critical assumptions by requesting additional evidence or verification from subordinate agents. 3. Log meta-data about agent reliability and final decision paths for interpretability. Example Usage: Conflict Re-emerges: Repeated contradictory statements about a single piece of evidence. Your Action: Issue a "challenge" directive to the Verifier Agent or the Retriever Agent, requesting additional sources or alternative cross-checks. Output Format (JSON Only): { "intervention_type": "challenge | override | escalate | none", "target_of_intervention": "Which agent or assertion is challenged", "rationale": "Explanation of why the Controller intervened", "meta_notes": "Optional additional commentary or reliability signals" } Instruction: Produce valid JSON only. This agent acts rarely, but can do so when repeated failures occur or when a major conflict must be forcibly resolved.
A.7 Usage and Integration Notes
The above prompts constitute a cooperative multi-agent system designed for reversible multi-hop question answering. The usage scenario is as follows:
  1. 1.
    Question-Decomposer AgentA.1) receives the user’s original query and splits it into sub-questions.
  2. 2.
    Retriever AgentA.2) fetches relevant information for each sub-question from external sources (text or KG).
  3. 3.
    Verifier AgentA.3) checks for local inconsistencies and can trigger local backtracking if contradictory evidence arises.
  4. 4.
    Answer-Assembler AgentA.4) merges partial answers into a final solution. If irreconcilable conflicts appear, it escalates to the Supervisor.
  5. 5.
    Supervisor AgentA.5) coordinates global resolution or wide-scale backtracking if multiple agents’ statements are in conflict.
  6. 6.
    Controller AgentA.6) optionally intervenes if repeated or severe conflicts persist, forcing extra checks or overrides.
This architecture supports non-monotonic reasoning, allowing the system to roll back to earlier states to correct errors and ensure robust, interpretable multi-hop QA. “Local backtracking” is handled by individual agents (especially the Verifier), whereas “global backtracking” is orchestrated by the Supervisor Agent when local corrections are insufficient.
All final outputs from each agent must adhere to the specified JSON schema to ensure interoperability. Downstream agents read the previous agent’s JSON fields directly, enabling a structured, chain-of-thought style pipeline that remains reversible at every step.
Appendix B Case Study
B.1 Puzzle Statement
Suspects and Rules:
  • There are four suspects: A, B, C, D.
  • Exactly one of them is the culprit.
  • The culprit must have at least one false statement. Non-culprits may also have errors, but are not forced to be entirely truthful or untruthful.
  • Each suspect makes the following statements:
    1. 1.
      A:
      1. (a)
        “I did not do it.”
      2. (b)
        “If B is the culprit, then C is lying (i.e., at least one statement from C is false).”
    2. 2.
      B:
      1. (a)
        “Either A is lying, or D is the culprit.”
      2. (b)
        (No second statement given in some puzzle variants, or it might be omitted. We assume just one statement from B here.)
    3. 3.
      C:
      1. (a)
        “B did not do it.”
      2. (b)
        “D has at least one untrue statement.”
    4. 4.
      D:
      1. (a)
        “C is lying about everything.” (i.e., both C(i) and C(ii) are false)
      2. (b)
        “I definitely did not do it.”
Question:
Which single suspect is guilty, respecting the puzzle rules?
B.2 Agents and Their Roles
  1. 1.
    Question-Decomposer Agent ($A_{Q}$)
    Splits the complex question into sub-questions:
    • Sub-question 1: Which U.S. state hosted the 1984 Summer Olympics?
    • Sub-question 2: What is the capital city of that state, and what is the largest city?
    • Sub-question 3: Compare the populations of the capital and the largest city.
    • Sub-question 4: Return the state name if the capital’s population is smaller.
  2. 2.
    Retriever Agent ($A_{R}$)
    Searches external knowledge (e.g., a text corpus or knowledge graph) to gather relevant facts:
    • Text passages or data about U.S. states, capitals, largest cities, and historical Olympic hosts.
  3. 3.
    Verifier Agent ($A_{V}$)
    Cross-checks newly retrieved information for consistency against its local knowledge or prior verified facts. If a contradiction is detected, it triggers local backtracking.
  4. 4.
    Answer-Assembler Agent ($A_{A}$)
    Synthesizes partial answers from the other agents. If contradictory partial answers cannot be reconciled at the local level, $A_{A}$ escalates the conflict to the Supervisory layer.
  5. 5.
    Supervisor Agent ($A_{S}$)
    Oversees system-wide conflicts. It can perform global backtracking (rolling back all agents) if needed.
  6. 6.
    Controller Agent ($A_{C}$)
    Provides strategic oversight. In the example below, it will issue “challenges” to specific assertions when local backtracking fails.
B.3 Walkthrough of the Reasoning and Backtracking
B.3.1 Initial Question Decomposition
  1. 1.
    $A_{Q}$ receives the question:
    $A_{Q}$ parses the main query:
    “Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?”
  2. 2.
    $A_{Q}$ produces sub-questions:
    • $q_{1}$: “Which U.S. state hosted the 1984 Summer Olympics?”
    • $q_{2}$: “What is the capital city of that state? Also, what is the largest city of that state?”
    • $q_{3}$: “Compare the population of the capital city to that of the largest city.”
    • $q_{4}$: “Return the state if the capital’s population is smaller.”
    These sub-questions ($q_{1}$$q_{4}$) are broadcast to the Retriever ($A_{R}$) and Verifier ($A_{V}$).
B.3.2 Retrieval and a Local Conflict
  1. 1.
    $A_{R}$ retrieves answers:
    • For $q_{1}$, $A_{R}$ returns:
      “The 1984 Summer Olympics were primarily hosted in Los Angeles, California.”
    • For $q_{2}$, $A_{R}$ retrieves partial data:
      • Capital of California is Sacramento.
      • Largest city is Los Angeles.
      • Population estimates (unfortunately, one item is erroneous or inconsistent):
          Sacramento population: $\sim 508,000$ (from official records).
          Los Angeles population: $\sim 3,900,000$ (correct).
          But a second retrieved record incorrectly lists Sacramento’s population as 1,500,000.
  2. 2.
    $A_{V}$ (Verifier) checks local consistency:
    $A_{V}$ notices that two different population figures for Sacramento are contradictory:
    • Data point $D_{1}$: 508,000
    • Data point $D_{2}$: 1,500,000
    $A_{V}$ identifies these as mutually exclusive facts. It flags a local conflict:
    “Population of Sacramento cannot be both 508k and 1.5M.”
  3. 3.
    Local Backtracking by $A_{V}$:
    Before finalizing any partial answer, $A_{V}$ backtracks to the checkpoint prior to adopting the second, suspicious data point $D_{2}$.
    $A_{V}$ discards the contradictory population figure (1.5M) and retains only 508k as the consistent number for Sacramento.
    The system continues forward with the corrected value for Sacramento’s population: $\sim 508,000$.
(This demonstrates how a single agent can retract contradictory or low-confidence evidence without halting the entire process.)
B.3.3 Assembling Partial Answers & Uncovering a Global Conflict
  1. 1.
    $A_{A}$ (Answer-Assembler) integrates partial conclusions:
    So far, the state that hosted the 1984 Olympics is California.
    The capital is Sacramento (population: $\sim 508,000$).
    The largest city is Los Angeles (population: $\sim 3,900,000$).
  2. 2.
    $A_{A}$ composes a preliminary final:
    Since Sacramento’s population ($\sim 508,000$) is indeed smaller than Los Angeles’s ($\sim 3,900,000$), the preliminary answer is: “California” should be the correct state.
  3. 3.
    New conflict introduced:
    However, suppose the Retriever ($A_{R}$)—in parallel—fetched an alternative “capital city” record stating that Los Angeles was once referred to as the “capital” in a historical context (erroneous snippet from a non-authoritative source).
    This implies the contradictory statement: “Los Angeles is also the capital of California,” which directly conflicts with the known fact “Sacramento is the capital of California.”
  4. 4.
    $A_{V}$ tries local resolution:
    $A_{V}$ cannot reconcile “Los Angeles is capital” with “Sacramento is capital.”
    Because each piece of evidence was introduced into separate sub-threads, a single local reversion inside $A_{V}$ might not suffice. The conflict is also recognized by $A_{A}$ when finalizing sub-answers.
  5. 5.
    Conflict escalates to the Supervisor ($A_{S}$):
    As soon as multiple agents disagree over fundamental facts (the identity of the capital), the system triggers a global conflict signal.
    $A_{S}$ identifies the minimal conflicting set:
    • Capital(California, Sacramento)
    • Capital(California, Los Angeles)
    These two are clearly incompatible. The next step is to identify which statement should be retracted system-wide.
B.3.4 Global Backtracking and Final Resolution
  1. 1.
    Global Backtracking:
    The Supervisor Agent ($A_{S}$) issues a global backtracking command, rolling the entire system’s knowledge to a shared checkpoint before the contradictory capital reference was accepted.
    All local knowledge caches revert to a consistent state in which “Sacramento is the capital” is still true, and “Los Angeles is the capital” is no longer present.
  2. 2.
    Controller ($A_{C}$) challenges the suspicious assertion:
    To prevent the same contradiction from reappearing, $A_{C}$ explicitly “challenges” the statement Capital(California, Los Angeles).
    This statement is reevaluated or ignored based on domain knowledge or reliability checks (e.g., confidence weighting from the retriever and the verifier).
    The system confirms that Los Angeles was never the official capital.
  3. 3.
    Assemble the final consistent answer:
    With the conflicting statement removed, the pipeline reaffirms:
    1. (i)
      The 1984 Summer Olympics took place in Los Angeles, California.
    2. (ii)
      California’s capital is Sacramento.
    3. (iii)
      Sacramento’s population ($\sim 508,000$) is smaller than Los Angeles’s ($\sim 3,900,000$).
    Final Answer: California.
B.4 Key Observations
  • Early-Stage Conflict Resolution: The Verifier Agent ($A_{V}$) performed local backtracking when it discovered contradictory population data for Sacramento. This promptly removed an incorrect data point without halting the entire process.
  • Escalation of Irreconcilable Contradictions: When two different sub-threads provided fundamentally clashing information (competing claims about the capital), the system automatically escalated the conflict to the Supervisor ($A_{S}$) for global action.
  • Strategic Re-check and Override: The Controller Agent ($A_{C}$) could forcibly challenge the suspicious statement “Capital(California, Los Angeles)” to eliminate it from the knowledge pool, thus preserving the correct solution.
Overall, this case exemplifies how reversible, multi-hop reasoning allows for robust error correction: local invalid data is undone swiftly, while deeper logic conflicts trigger system-wide backtracking. Consequently, the answer “California” remains stable and correct, ensuring that a single faulty retrieval step does not irreversibly corrupt the entire reasoning chain.
Appendix C Traditional Multi-hop QA Tasks
1. The Question
C.0.1 User’s question:
"Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?"
To answer this, the system must:
  1. 1.
    Identify the state that hosted the 1984 Summer Olympics.
  2. 2.
    Compare the capital city’s population to the largest city’s population in that state.
  3. 3.
    Confirm the capital city’s population is indeed smaller.
  4. 4.
    Provide the name of that state.
Although it may appear straightforward, we deliberately introduce contradictory or erroneous data along the way to showcase the reversible (backtracking) mechanisms.
C.0.2 Agents and Their Roles
  1. 1.
    Question-Decomposer Agent ($A_{Q}$)
    Splits the complex question into sub-questions:
    • Sub-question 1: Which U.S. state hosted the 1984 Summer Olympics?
    • Sub-question 2: What is the capital city of that state, and what is the largest city?
    • Sub-question 3: Compare the populations of the capital and the largest city.
    • Sub-question 4: Return the state name if the capital’s population is smaller.
  2. 2.
    Retriever Agent ($A_{R}$)
    Searches external knowledge (e.g., a text corpus or knowledge graph) to gather relevant facts:
    • Text passages or data about U.S. states, capitals, largest cities, and historical Olympic hosts.
  3. 3.
    Verifier Agent ($A_{V}$)
    Cross-checks newly retrieved information for consistency against its local knowledge or prior verified facts. If a contradiction is detected, it triggers local backtracking.
  4. 4.
    Answer-Assembler Agent ($A_{A}$)
    Synthesizes partial answers from the other agents. If contradictory partial answers cannot be reconciled at the local level, $A_{A}$ escalates the conflict to the Supervisory layer.
  5. 5.
    Supervisor Agent ($A_{S}$)
    Oversees system-wide conflicts. It can perform global backtracking (rolling back all agents) if needed.
  6. 6.
    Controller Agent ($A_{C}$)
    Provides strategic oversight. In the example below, it will issue “challenges” to specific assertions when local backtracking fails.
C.0.3 Walkthrough of the Reasoning and Backtracking
3.1 Initial Question Decomposition
  1. 1.
    $A_{Q}$ receives the question:
    $A_{Q}$ parses the main query:
    “Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?”
  2. 2.
    $A_{Q}$ produces sub-questions:
    • $q_{1}$: “Which U.S. state hosted the 1984 Summer Olympics?”
    • $q_{2}$: “What is the capital city of that state? Also, what is the largest city of that state?”
    • $q_{3}$: “Compare the population of the capital city to that of the largest city.”
    • $q_{4}$: “Return the state if the capital’s population is smaller.”
    These sub-questions ($q_{1}$$q_{4}$) are broadcast to the Retriever ($A_{R}$) and Verifier ($A_{V}$).
3.2 Retrieval and a Local Conflict
  1. 1.
    $A_{R}$ retrieves answers:
    • For $q_{1}$, $A_{R}$ returns:
      “The 1984 Summer Olympics were primarily hosted in Los Angeles, California.”
    • For $q_{2}$, $A_{R}$ retrieves partial data:
      • Capital of California is Sacramento.
      • Largest city is Los Angeles.
      • Population estimates (unfortunately, one item is erroneous or inconsistent):
          Sacramento population: $\sim 508,000$ (from official records).
          Los Angeles population: $\sim 3,900,000$ (correct).
          But a second retrieved record incorrectly lists Sacramento’s population as 1,500,000.
  2. 2.
    $A_{V}$ (Verifier) checks local consistency:
    $A_{V}$ notices that two different population figures for Sacramento are contradictory:
    • Data point $D_{1}$: 508,000
    • Data point $D_{2}$: 1,500,000
    $A_{V}$ identifies these as mutually exclusive facts. It flags a local conflict:
    “Population of Sacramento cannot be both 508k and 1.5M.”
  3. 3.
    Local Backtracking by $A_{V}$:
    Before finalizing any partial answer, $A_{V}$ backtracks to the checkpoint prior to adopting the second, suspicious data point $D_{2}$.
    $A_{V}$ discards the contradictory population figure (1.5M) and retains only 508k as the consistent number for Sacramento.
    The system continues forward with the corrected value for Sacramento’s population: $\sim 508,000$.
(This demonstrates how a single agent can retract contradictory or low-confidence evidence without halting the entire process.)
3.3 Assembling Partial Answers & Uncovering a Global Conflict
  1. 1.
    $A_{A}$ (Answer-Assembler) integrates partial conclusions:
    So far, the state that hosted the 1984 Olympics is California.
    The capital is Sacramento (population: $\sim 508,000$).
    The largest city is Los Angeles (population: $\sim 3,900,000$).
  2. 2.
    $A_{A}$ composes a preliminary final:
    Since Sacramento’s population ($\sim 508,000$) is indeed smaller than Los Angeles’s ($\sim 3,900,000$), the preliminary answer is: “California” should be the correct state.
  3. 3.
    New conflict introduced:
    However, suppose the Retriever ($A_{R}$)—in parallel—fetched an alternative “capital city” record stating that Los Angeles was once referred to as the “capital” in a historical context (erroneous snippet from a non-authoritative source).
    This implies the contradictory statement: “Los Angeles is also the capital of California,” which directly conflicts with the known fact “Sacramento is the capital of California.”
  4. 4.
    $A_{V}$ tries local resolution:
    $A_{V}$ cannot reconcile “Los Angeles is capital” with “Sacramento is capital.”
    Because each piece of evidence was introduced into separate sub-threads, a single local reversion inside $A_{V}$ might not suffice. The conflict is also recognized by $A_{A}$ when finalizing sub-answers.
  5. 5.
    Conflict escalates to the Supervisor ($A_{S}$):
    As soon as multiple agents disagree over fundamental facts (the identity of the capital), the system triggers a global conflict signal.
    $A_{S}$ identifies the minimal conflicting set:
    • Capital(California, Sacramento)
    • Capital(California, Los Angeles)
    These two are clearly incompatible. The next step is to identify which statement should be retracted system-wide.
3.4 Global Backtracking and Final Resolution
  1. 1.
    Global Backtracking:
    The Supervisor Agent ($A_{S}$) issues a global backtracking command, rolling the entire system’s knowledge to a shared checkpoint before the contradictory capital reference was accepted.
    All local knowledge caches revert to a consistent state in which “Sacramento is the capital” is still true, and “Los Angeles is the capital” is no longer present.
  2. 2.
    Controller ($A_{C}$) challenges the suspicious assertion:
    To prevent the same contradiction from reappearing, $A_{C}$ explicitly “challenges” the statement Capital(California, Los Angeles).
    This statement is reevaluated or ignored based on domain knowledge or reliability checks (e.g., confidence weighting from the retriever and the verifier).
    The system confirms that Los Angeles was never the official capital.
  3. 3.
    Assemble the final consistent answer:
    With the conflicting statement removed, the pipeline reaffirms:
    1. (i)
      The 1984 Summer Olympics took place in Los Angeles, California.
    2. (ii)
      California’s capital is Sacramento.
    3. (iii)
      Sacramento’s population ($\sim 508,000$) is smaller than Los Angeles’s ($\sim 3,900,000$).
    Final Answer: California.
C.0.4 4. Key Observations
  • Early-Stage Conflict Resolution: The Verifier Agent ($A_{V}$) performed local backtracking when it discovered contradictory population data for Sacramento. This promptly removed an incorrect data point without halting the entire process.
  • Escalation of Irreconcilable Contradictions: When two different sub-threads provided fundamentally clashing information (competing claims about the capital), the system automatically escalated the conflict to the Supervisor ($A_{S}$) for global action.
  • Strategic Re-check and Override: The Controller Agent ($A_{C}$) could forcibly challenge the suspicious statement “Capital(California, Los Angeles)” to eliminate it from the knowledge pool, thus preserving the correct solution.
Overall, this case exemplifies how reversible, multi-hop reasoning allows for robust error correction: local invalid data is undone swiftly, while deeper logic conflicts trigger system-wide backtracking. Consequently, the answer “California” remains stable and correct, ensuring that a single faulty retrieval step does not irreversibly corrupt the entire reasoning chain.

### v2

@@
 Abstract
-Recent advances in large language models (LLMs) have significantly improved multi-hop question answering (QA) through direct Chain-of-Thought (CoT) reasoning.
-However, the irreversible nature of CoT leads to error accumulation, making it challenging to correct mistakes in multi-hop reasoning.
-This paper introduces ReAgent: a Reversible multi-Agent collaborative framework augmented with explicit backtracking mechanisms, enabling reversible multi-hop reasoning.
-By incorporating text-based retrieval, information aggregation and validation, our system can detect and correct errors mid-reasoning, leading to more robust and interpretable QA outcomes.
-The framework and experiments serve as a foundation for future work on error-tolerant QA systems.
-Empirical evaluations across three benchmarks indicate ReAgent’s efficacy, yielding average about 6% improvements against baseline models.
-ReAgent: Reversible Multi-Agent Reasoning for Knowledge-Enhanced Multi-Hop QA
-Zhao Xinjie1, Fan Gao1, Rui Yang2, Yingjian Chen1, Yuyang Wang3, Ying Zhu1, Jiacheng Tang1, Irene Li1 1 The University of Tokyo, 2 Duke-NUS Medical School, 3 The Hong Kong University of Science and Technology (Guangzhou)
+Multi-hop question answering (QA) remains challenging, as solutions must reliably integrate and reconcile evidence from multiple sources without succumbing to error propagation.
+While large language models (LLMs) have achieved substantial improvements via chain-of-thought (CoT) prompting and retrieval-augmented generation, these methods typically adopt a forward-only workflow—early mistakes persist throughout inference, and contradictions discovered later cannot systematically trigger re-evaluation.
+To address this limitation, we present ReAgent, a reversible multi-agent reasoning framework.
+Specifically, ReAgent enables agents to backtrack to earlier valid states when conflicts arise, thereby isolating and rectifying flawed assumptions before they undermine subsequent reasoning.
+Our approach combines explicit local and global rollback protocols with modular role specialization, resulting in a flexible and error-tolerant pipeline.
+Empirical evaluation on three multi-hop QA benchmarks demonstrates consistent performance gains of approximately 6% over forward-only baselines, in addition to enhanced interpretability.
+These findings highlight the value of non-monotonic, backtracking-driven inference in complex QA scenarios and point to broader implications for multi-agent collaboration in knowledge-intensive tasks.
+[^1] Our anonymous code is available at https://anonymous.4open.science/r/ReAgent-9415.
+ReAgent: Reversible Multi-Agent Reasoning for
+Knowledge-Enhanced Multi-Hop QA
+Xinjie Zhao1$\dagger$, Fan Gao1$\dagger$, Xingyu Song1, Yingjian Chen1, Rui Yang2, Yanran Fu1, Yuyang Wang3, Yusuke Iwasawa1, Yutaka Matsuo1, Irene Li1 1The University of Tokyo 2Duke–NUS Medical School 3The Hong Kong University of Science and Technology irene.li@weblab.t.u-tokyo.ac.jp
+$\dagger$\$$\dagger$\$footnotetext: Equal contribution.
 1 Introduction
-Multi-hop question answering (QA) represents a crucial challenge in natural language processing (NLP), requiring systems to integrate and reconcile evidence from multiple sources or hops of reasoning before reaching a single final answer (Yang et al., 2018; Welbl et al., 2018; Ho et al., 2020).
-Existing benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020) have shed light on the intrinsic difficulty of multi-hop inference: not only must relevant information be located across dispersed documents or knowledge graphs, but the process itself must remain robust to partial mistakes.
-If an early inference is incorrect, subsequent steps can amplify that error and produce an unreliable or contradictory final answer (Inoue et al., 2020; Bo et al., 2024; Yang et al., 2024c).
-Recent large language models (LLMs) have showcased strong performance on multi-hop QA, often employing either explicit Chain-of-Thought (CoT) prompting (Wang et al., 2023) or retrieval-augmented generation (Das et al., 2018; Long et al., 2019; Gao et al., 2023; Liu et al., 2023; Yang et al., 2024b, 2025).
-While these strategies can help track intermediate steps, they typically rely on a predominantly forward-directed workflow.
-Once a partial conclusion is formed, the system rarely revisits or corrects it unless externally prompted (Puerto et al., 2023; Huang et al., 2024).
-This forward-only structure poses limitations, especially when later evidence contradicts an earlier inference or when multiple lines of reasoning yield divergent findings.
-A single incorrect hop can invalidate the final answer, and without explicit backtracking, the system is unable to retract and repair flawed assumptions (Doyle, 1979; Bo et al., 2024).
-Figure 1: We compare different frameworks for QA: Linear (CoT), Collaborative (MAS), and Reversible (ReAgent) Reasoning in Multi-hop QA.
-To address this gap, growing attention has been directed to multi-agent collaboration for QA, where specialized agents focus on sub-tasks—such as retrieval, verification, and resolution of conflicting information—and then synchronize their outcomes (Zhao et al., 2024; Parhizkar et al., 2020; Ke et al., 2024).
-By delegating different roles (e.g., question decomposition, result verification) to different agents, the system becomes more modular and potentially more explainable.
-Recent multi-agent systems include COPPER (Bo et al., 2024) and LongAgent (Zhao et al., 2024), both of which decompose complex tasks among multiple LLM-based agents that coordinate via message passing.
-They not only reduce overall context per agent but can also cross-check for consistency, flagging suspicious partial solutions.
-However, even these advanced frameworks often lack fully reversible reasoning pipelines, wherein each agent can effectively backtrack to an earlier valid state once a conflict is detected, re-explore alternatives, and propagate the corrections forward (He et al., 2021).
-In this paper, we introduce ReAgent, a reversible multi-agent collaborative reasoning framework with explicit backtracking, designed for multi-hop question answering.
-This approach builds on the premise that in real-world complex QA scenarios—especially when reasoning across a broad knowledge graph or multiple textual sources—one-size-fits-all steps fail to capture the frequent need for partial re-evaluation (Yang et al., 2018; Ho et al., 2020).
-Conflict detection and backtracking mechanisms act as a flexible error-correction loop, in which a mismatch in evidence can trigger a re-check of previously accepted facts, ensuring that misinterpretations do not survive to the final answer.
-This perspective aligns with “reflective” multi-agent designs (Bo et al., 2024; Huang et al., 2024), but pushes further by defining explicit multi-layer protocols—local versus global backtracking—and employing a dynamic aggregator that leverages trust signals from each agent’s past reliability (Puerto et al., 2023; Parhizkar et al., 2020).
-We empirically show that this reversible approach improves final-answer accuracy and interpretability: backtracking traces clarify how the system isolates and corrects mistakes.
-Our work links robust non-monotonic reasoning with data-driven QA, offering a transparent, error-tolerant multi-hop inference pipeline.
-In summary, our main contributions are: (1) a novel multi-agent QA framework that supports backtracking at both local and global levels; (2) hybrid retrieval from knowledge graphs and text for flexible evidence acquisition; and (3) systematic experiments on three benchmark datasets demonstrating more robust and interpretable multi-hop QA than forward-only baselines.
-Specifically, our framework achieves an average improvement of approximately 6% over baseline models.
-Figure 2: Execution, Supervision, and Interaction Layers: Three-Fold Role Partitioning in ReAgent’s Multi-Hop QA Backtracking Mechanisms.
+Multi-hop question answering (QA) is a central challenge in natural language processing (NLP), demanding the ability to gather and integrate evidence across multiple documents, database entries, or knowledge-graph nodes before converging on a single correct answer (Yang et al., 2018; Welbl et al., 2018; Ho et al., 2020).
+Benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020) highlight the intricate nature of multi-hop inference, where each reasoning step can involve partial retrieval, validation, and synthesis of new information.
+A core difficulty lies in the system’s vulnerability to early mistakes: if an incorrect inference is made at an initial hop, subsequent steps often propagate this error and undermine the final outcome, rendering it contradictory or simply wrong (Inoue et al., 2020; Bo et al., 2024; Yang et al., 2024c).
+This phenomenon has motivated extensive research into chaining intermediate inferences and exploring ways to detect, isolate, or rectify problematic conclusions.
+Figure 1: Comparison of multi-hop reasoning strategies.
+Chain-of-Thought (CoT) and Multi-Agent Systems (MAS) typically adopt a forward-driven reasoning pipeline without rollback mechanisms, which could generate the wrong answer due to error accumulation.
+In contrast, our proposed ReAgent introduces explicit backtracking mechanisms that enable the system to correct errors during reasoning, resulting in a more accurate and reliable answer.
+Recent large language models (LLMs) exhibit promising results on multi-hop QA, frequently using either explicit Chain-of-Thought (CoT) prompting (Wang et al., 2023) (Figure 1, left) or retrieval-augmented generation (Das et al., 2018; Long et al., 2019; Gao et al., 2023; Liu et al., 2023; Yang et al., 2024b, 2025a).
+These methods facilitate a stepwise approach, encouraging transparency in how each intermediate fact is reached.
+Nonetheless, they typically rely on a forward-driven reasoning pipeline that does not proactively examine or correct previously accepted statements.
+In practice, once a model commits an erroneous partial inference may not reevaluate it unless given targeted prompts to do so (Puerto et al., 2023; Huang et al., 2024).
+This unidirectional paradigm is problematic when later-discovered evidence or reasoning paths contradict prior assumptions, as the system lacks a structured mechanism to revise and propagate corrections.
+Although incremental improvements have been proposed, the absence of robust backtracking or rollback still limits their reliability and interpretability (Doyle, 1979; Bo et al., 2024).
+A growing body of work in multi-agent collaboration (Figure 1, middle) offers an alternative perspective, assigning specialized components distinct roles in retrieval, validation, conflict detection, and results assembly (Zhao et al., 2024; Parhizkar et al., 2020; Ke et al., 2024).
+Approaches such as COPPER (Bo et al., 2024) and LongAgent (Zhao et al., 2024) distribute the QA task among multiple LLM-based agents that communicate via message passing, thereby providing greater modularity and a clearer division of labor.
+By cross-verifying evidence, each agent can potentially identify suspicious partial solutions.
+However, even these designs often lack a systematic strategy to revert to an earlier, valid state once a global contradiction emerges.
+Such a reversal capability is non-trivial, as it introduces synchronization complexities among the agents and raises questions about how to detect, prioritize, and handle contradictory or low-confidence information in a large-scale collaborative setting (He et al., 2021).
+In this paper, we propose ReAgent, a reversible multi-agent collaborative reasoning framework for multi-hop QA (Figure 1, right).
+Our method introduces explicit backtracking protocols into a multi-layer architecture that addresses both the granular aspects of local error correction and the broader system-wide consistency checks.
+ReAgent alleviates error accumulation in forward-only strategies by incorporating fine-grained conflict-detection cues at each step and using a flexible error-correction loop that can revert and iteratively refine intermediate inferences.
+Our design is inspired by multi-agent systems with reflective capabilities (Bo et al., 2024; Huang et al., 2024), but we go further by defining how local versus global backtracking unfolds, how to manage concurrency issues when parallel agents must revert their states, and how to integrate trust signals derived from each agent’s past reliability (Puerto et al., 2023; Parhizkar et al., 2020).
+While the addition of reversible reasoning raises legitimate questions about computational overhead and concurrency, our findings show that these challenges can be mitigated by careful coordination at the supervisory level.
+Experiments on three public multi-hop QA benchmarks demonstrate that ReAgent improves final-answer accuracy while enhancing interpretability, and even outperforms several advanced reasoning models despite it built on a lightweight, non-reasoning foundation.
+Figure 2: The overall architecture of the ReAgent.
+The given question is processed through the Execution Layer, which involves question decomposition, evidence retrieval, verification, and is ultimately integrated to generate the final answer (blue line).
+The Supervisor Layer and Interaction Layer are responsible for monitoring, regulation, and communication.
+The ReAgent framework includes both local and global backtracking mechanisms (red boxes), triggered by the Verifier Agent ($A_{V}$) and Supervisor Agent ($A_{S}$), respectively.
+Our contributions are threefold: (1) we propose a multi-agent QA framework that supports both local and global backtracking to correct mistakes in situ; (2) we design a hybrid retrieval mechanism that integrates textual and graph-based evidence; and (3) We empirically demonstrate the effectiveness of ReAgent, achieving an average accuracy gain of approximately 6% over the strongest baselines, while improving robustness and transparency compared to forward-only methods.
+2 Related Work
+Prompting and Iterative Reasoning.
+Large Language Models (LLMs) can tackle complex tasks using chain-of-thought (CoT) promptingYang et al.
+(2024a, 2025b), which promotes step-by-step reasoning and improves performance in arithmetic, commonsense, and symbolic tasks (Wei et al., 2022).
+Accuracy further improves with self-consistency, which samples multiple reasoning paths and selects the majority answer (Wang et al., 2023).
+To reduce manual prompt design, automatic methods such as APE generate and evaluate prompts via LLMs themselves, treating prompt creation as a search task (Zhou et al., 2023).
+Iterative frameworks like ReAct combine reasoning with tool use to refine answers (Yao et al., 2023), while Reflexion and Self-Refine add feedback loops where the model critiques and revises its outputs (Shinn et al., 2023; Madaan et al., 2023).
+Prompting has also extended to multimodal inputs; for instance, Multimodal-CoT integrates visual and textual information into a joint reasoning trace to generate the final answer (Zhang et al., 2023).
+Multi-Agent Collaboration, Debate, and Scalability.
+Recent work investigates multi-agent systems where multiple LLMs collaborate or compete toward shared goals.
+CAMEL assigns roles (e.g., “user” and “assistant”) to agents that communicate via dialogue to decompose tasks (Li et al., 2023).
+Role specialization and consensus mechanisms enhance robustness.
+Debate-based methods push this further, with agents arguing opposing views and a judge—human or model—selecting the best solution (Khan et al., 2024; Xiong et al., 2023).
+Such adversarial exchanges promote correctness, though dominant agents can bias results without capable judges.
+As agent counts grow, coordination becomes a bottleneck.
+Efficient structures like hierarchical controllers and learned protocols are needed to manage scalability and communication overhead (Li et al., 2023; Xiong et al., 2023).
 3 Preliminary
 […]
-Challenges in Forward-Only Reasoning A key source of error in multi-hop QA lies in the lack of revision capacity.
-Once a system commits to a specific inference early on (for instance, deducing a partial conclusion from $e_{1}$ and $e_{2}$), there is no built-in mechanism to retract or update that inference if additional evidence suggests it is incorrect.
-As the system progresses to hop $j{+}1$ or further, it carries potential inaccuracies along, making the final answer unreliable or inconsistent.
-In many cases, the model would have benefited from re-examining initial assumptions, especially when encountering contradictory or low-confidence signals.
-These observations motivate an interactive and flexible approach, wherein the QA process is allowed to backtrack and replace problematic intermediate results.
-Non-Monotonic Backtracking Our notion of backtracking is a non-monotonic extension of the typical multi-hop QA process.
-Rather than treating the knowledge state as strictly accumulative, we allow for partial or global reversion.
-Specifically:
-Local Correction: An agent can retract one of its own inferences if it detects internal conflicts or if it receives contradictory evidence from other agents.
-Global Rollback: When inconsistencies span multiple agents, a supervisory mechanism coordinates a synchronized rollback.
-This can mean discarding a sequence of accepted statements that are ultimately untenable, returning the system to a previously consistent state.
-This extended capability is guided by two key considerations.
-First, not all new evidence is guaranteed to be more credible than old; hence, the system must weigh or prioritize contradictory claims.
-Second, if an inference $I$ made in hop $r$ is discredited in hop $r{+}2$, it may be more efficient to rollback to hop $r$ directly, rather than forcing the pipeline to proceed with a known contradiction.
-Such rollback actions effectively prevent one faulty hop from invalidating the entire chain of reasoning.
-Figure 3: Layered Multi-Agent Interaction and Backtracking Flow with Time Step.
+Non-Monotonic Backtracking.
+Our notion of backtracking is a non-monotonic extension of the typical multi-hop QA process.
+We introduce a reversible reasoning mechanism that allows both local and global backtracking.
+  • Local Correction: An agent can revise its own inference when it detects internal conflict or receives contradictory evidence from other agents.
+  • Global Rollback: A supervisor coordinates rollback across agents when inconsistencies span multiple agents, restoring the system to a previously consistent state.
 4 ReAgent: Reversible Multi-Agent Reasoning Architecture
-3.1 Overview of the ReAgent Architecture
-Figure 2 presents the overall design, organized into three layers.
-Let $\mathcal{A}\;=\;\{A_{1},A_{2},\dots,A_{n}\}$ be the set of agents.
-At time step $t$, each agent $A_{i}$ maintains a local knowledge set $\Phi_{i}^{t}$ that may include proposed assertions, retrieved evidence, or partial inferences.
-In addition, the system can track a global assertion set $\Phi^{t}\;=\;\bigcup_{i=1}^{n}\Phi_{i}^{t},$ representing all statements under consideration at time $t$.
-The three-layer structure is as follows:
-  1. 1.
-    Execution Layer (Section 3.2): Comprises the essential agents for domain tasks, such as decomposing questions, retrieving evidence, verifying local consistency, and assembling intermediate answers.
-  2. 2.
-    Supervisory Layer (Section 3.3): Includes supervisory agents that orchestrate conflict resolution and manage global backtracking, ensuring that contradictory or invalid assertions are addressed across all relevant agents.
-  3. 3.
-    Interaction Layer (Section 3.4): Maintains the concurrency model and communication protocols.
-It defines how agents send and receive messages (e.g., assert, reject, challenge) and how the system merges these parallel events into consistent updates of $\Phi^{t}$.
-By separating concerns across these layers, it becomes simpler to reason about local agent updates while still allowing global interventions.
+Figure 2 presents the overall architecture of our proposed ReAgent, organized into three layers: 1) Execution Layer, responsible for decomposing the input question $Q$ into multiple sub-questions, retrieving relevant evidence respectively, validating intermediate results, and integrating them to generate the final answer $A_{final}$; 2) Supervisor Layer, responsible for high-level regulation, coordinating conflict resolution and managing global backtracking; and 3) Interaction Layer, responsible for maintaining the concurrency model and communication protocols.
+The core of the architecture is a hierarchical backtracking mechanism, consisting of local backtracking, which resolves internal contradictions within each agent, and global backtracking, which handles contradictions spanning multiple agents.
+To support this, ReAgent maintains knowledge sets at each time step $t$.
+Specifically, given a set of agents $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$, each agent $A_{i}$ holds a local knowledge set $\Phi_{i}^{t}$, containing its proposed assertions, retrieved evidence, or intermediate inferences, while the system maintains a global knowledge set $\Phi^{t}=\bigcup_{i=1}^{n}\Phi_{i}^{t}$, representing the set of global statements under consideration at time $t$.
 The design explicitly supports non-monotonic updates: newly introduced statements can be revoked if they lead to logical conflicts or are superseded by contradictory evidence.
-The specific prompts of each kind of agent is provided in Appendix A.
+The specific prompts of each kind of agent are provided in Appendix A.
 4.1 Execution Layer Agents
-The Execution Layer hosts four types of agents that address fundamental QA subtasks, each maintaining its local knowledge $\Phi_{i}$.
-When necessary, an agent may perform local backtracking to remove or adjust recent assertions.
+The Execution Layer hosts four types of agents that address fundamental QA sub-tasks, each maintaining its local knowledge $\Phi_{i}$.
 Question-Decomposer Agent ($A_{Q}$).
-Breaks a complex input question $Q_{0}$ into sub-questions $\{q_{1},q_{2},\dots\}$.
-If $Q_{0}(x)$ denotes the main query, $A_{Q}$ produces a decomposition mapping $\{q_{j}\}$ for subsequent retrieval and verification: $A_{Q}:\;Q_{0}(x)\;\longmapsto\;\{\,q_{1},\dots,q_{m}\}.$ This decomposition is broadcast to other base-layer agents.
+Given a complex input question $Q$, this agent breaks it into a set of sub-questions for subsequent retrieval and verification: $A_{Q}:Q\longmapsto\{q_{1},\dots,q_{m}\}$.
+The decomposition is broadcast to other base-layer agents.
 Retriever Agent ($A_{R}$).
-Upon receiving a sub-question $q_{i}$, $A_{R}$ returns relevant evidence $E$ (e.g., text passages or knowledge graph triples), possibly with confidence scores $\alpha(E)$.
-This agent may maintain a local retrieval cache and re-invoke searches if backtracking invalidates prior evidence.
+Upon receiving a sub-question $q_{i}$, $A_{R}$ issues parallel sparse and dense queries over the corpus, merges the hits with reciprocal-rank fusion, and retains the top-$M$ passages as the evidence set $E={e_{1},\dots,e_{M}}$.
+If backtracking invalidates any $e_{j}\in E$, only the associated $q_{i}$ is re-queried, leaving the rest unchanged.
 Verifier Agent ($A_{V}$).
 Performs local consistency checks.
-Given new evidence $E$, it verifies coherence with $\Phi_{V}^{t}$ (the Verifier’s current knowledge).
-If conflicts arise, $A_{V}$ triggers local backtracking to revert to a consistent state or signals higher-level agents for broader resolution.
-Concretely, $A_{V}$ may invoke
-$\mathrm{BacktrackLocal}\bigl{(}A_{V},r\bigr{)}$
-if it suspects that assertions introduced after node $r$ are inconsistent.
+Given a new evidence set $E$, it verifies coherence with its current knowledge $\Phi_{V}^{t}$.
+If conflicts arise, it triggers local backtracking to revert to a consistent state or signals higher-level agents for broader resolution.
+Concretely, $A_{V}$ may invoke $\mathrm{BacktrackLocal}\bigl{(}A_{V},r\bigr{)}$ if it detects inconsistencies in assertions introduced after node $r$.
 Answer-Assembler Agent ($A_{A}$).
-Gathers partial answers (and verified evidence) to synthesize a final solution.
-If the set of local inferences from $A_{Q}$, $A_{R}$, and $A_{V}$ is $\Phi_{\text{base}}$, then $A_{A}$ combines them to generate the final answer $A_{\text{final}}$.
-Any detected contradiction triggers further checks or possible escalation to the supervisory layer.
+Gathers partial answers (and verified evidence) to synthesize a final answer.
+Given the local inferences from $A_{Q}$, $A_{R}$, and $A_{V}$, the agent combines them to generate the final answer $A_{final}$, represented as:
+$A_{final}=A_{A}(\Phi_{Q},\Phi_{R},\Phi_{V}),$
+where $\Phi_{Q}$, $\Phi_{R}$ and $\Phi_{V}$ denotes the knowledge set from the respective agents. Any detected contradiction spanning multiple agents triggers escalation to the supervisory layer.
 4.2 Supervisory Layer Agents
-The supervisory layer manages conflicts that cannot be resolved by a single agent.
-It also oversees system-wide strategies, especially when contradictory goals or inconsistent states appear across multiple agents.
+The supervisory layer oversees system-wide strategies, especially when contradictory goals or inconsistent states appear across multiple agents that cannot be resolved by a single agent itself.
+Controller Agent ($A_{C}$).
+Regulates high-level strategies by monitoring game-theoretic signals or meta-rules.
+If a certain rule is deemed suboptimal or unsafe, $A_{C}$ can override it or enforce mode switches.
+For instance, it may issue a $\textsf{challenge}(\varphi)$ to examine a crucial assertion $\varphi$ from multiple agents’ perspectives or override local decisions if they jeopardize overall consistency.
 Supervisor Agent ($A_{S}$).
 Coordinates multi-agent conflicts spanning multiple knowledge sets or critical shared assumptions.
-If a conflict persists after local backtracking, $A_{S}$ determines whether partial or holistic rollback is required.
-When a system-wide contradiction is found, one may identify a minimal conflict set $\Psi$ such that $\mathrm{SAT}(\Psi)=\textsf{false}$ but $\mathrm{SAT}(\Psi^{\prime})=\textsf{true}$ for any proper subset $\Psi^{\prime}\subset\Psi$.
+If a conflict persists after local backtracking and escalation from $A_{A}$, the agent determines whether partial or holistic rollback is required.
+Specifically, when a system-wide contradiction is found, $A_{S}$ identifies a minimal conflict set $\Psi$ such that $\mathrm{SAT}(\Psi)=\textsf{false}$ but $\mathrm{SAT}(\Psi^{\prime})=\textsf{true}$ for any proper subset $\Psi^{\prime}\subset\Psi$.
 The supervisor then triggers a backtracking operation that eliminates or modifies $\Psi$.
-Controller Agent ($A_{C}$).
-Regulates high-level strategies, monitoring game-theoretic signals or meta-rules.
-If certain rules appear suboptimal or unsafe, $A_{C}$ can override them or enforce mode switches.
-For instance, it may issue a $\textsf{challenge}(\varphi)$ to examine a crucial assertion $\varphi$ from multiple agents’ perspectives, or override local decisions if they jeopardize overall consistency.
 4.3 Interaction Layer
-Communication Protocols.
-We adopt a message-driven framework, where atomic events $\mathtt{msg}(\varphi)$ are labeled with types such as assert, inform, reject, or challenge.
-These updates can occur simultaneously, requiring a composite event model to merge parallel messages.
-A shared concurrency mechanism processes simultaneous actions, ensuring that if conflicts arise from concurrent assertions, they are escalated to the supervisory layer.
-Temporal Operators.
-The system can utilize operators like $\square\varphi$ to denote that $\varphi$ must hold in every future state, and $\lozenge\varphi$ to indicate that a proposition $\varphi$ might become true at some future point.
+The Interaction Layer is responsible for storing knowledge sets from each interaction round, as well as maintaining the concurrency model and communication protocols.
+Persistent Log. Stores the local knowledge sets $\Phi_{i}$ and global knowledge set $\Phi$ from all interaction rounds to support backtracking and serve as a historical evidence repository.
+Temporal Tracker.
+Records the chronological sequence of messages and actions, enabling agents to reference previous steps accurately.
+Specifically, the system can use temporal operators such as $\square\varphi$ to denote that $\varphi$ must hold in every future state, and $\lozenge\varphi$ to indicate that a proposition $\varphi$ might become true at some future point.
 These temporal constraints assist in specifying persistent axioms or potential triggers for backtracking conditions.
-Algorithm 1 Multi-Agent Reversible Reasoning
-1:$Q_{0}$: Main question
-$A_{Q}.\mathrm{decompose}(Q_{0})\to\{q_{1},q_{2},\dots\}$
-3:Broadcast $\mathrm{assert}(q_{i})$ to $A_{R},A_{V},A_{A}$
-4:for each $q_{i}$ in $\{q_{1},q_{2},\dots\}$ concurrently do
-5:     $E\leftarrow A_{R}.\mathrm{retrieve}(q_{i})$
-6:     $A_{V}.\mathrm{verify}(E)$
-7:     if $A_{V}$ detects conflict locally then
-8:         $A_{V}.\mathrm{BacktrackLocal}(r_{V})$
-9:         if conflict persists then
-10:              raise Conflict to $A_{S}$
-11:         end if
-12:     end if
-13:     $A_{A}.\mathrm{storePartialAnswer}(q_{i},E)$
-14:end for
-15:if Global conflict signaled then
-16:     $A_{S}.\mathrm{HolisticUpdate}(\Pi_{j})$
-17:     if $A_{C}$ intervention needed then
-18:         $A_{C}.\mathrm{challenge}(\varphi)$ or override
-19:     end if
-20:end if
-21:Final $\leftarrow A_{A}.\mathrm{assembleAnswer}(\{q_{1},q_{2},\dots\})$
-22:return Final Answer
+Messaging Channel.
+Achieves communication across agents for exchanging updates, signaling conflicts, and broadcasting intermediate conclusions.
+Specifically, atomic events $\mathtt{msg}(\varphi)$ are typed as assert, inform, reject, or challenge.
+These updates can occur simultaneously, and a composite event model merges parallel messages.
+Furthermore, a shared concurrency mechanism processes simultaneous actions, ensuring that conflicts arising from concurrent assertions are escalated to the supervisory layer.
 4.4 Conflict Management and Backtracking
 This part formalizes how the system detects contradictions and reverts to consistent states.
-At each time $t$, the set of all accepted assertions is $\Phi^{t}\;=\;\bigcup_{i=1}^{n}\Phi_{i}^{t}.$ A conflict occurs if $\Phi^{t}$ entails both $\phi$ and $\neg\phi$ for some proposition $\phi$, meaning $\mathrm{SAT}(\Phi^{t})$ fails.
+At each time $t$, the set of all accepted assertions is $\Phi^{t}$.
+A conflict occurs if $\Phi^{t}$ entails both $\phi$ and $\neg\phi$ for some proposition $\phi$, meaning $\mathrm{SAT}(\Phi^{t})$ fails.
 To resolve the issue, the system follows a two-level approach:
@@
 Each agent $A_{i}$ maintains a backtracking graph $\mathrm{LBG}_{i}$, which logs states $\Phi_{i}^{r}$ at selected checkpoints $r$.
-If $\mathrm{Conflict}_{i}(\Phi_{i}^{t})$ indicates an internal contradiction, $A_{i}$ tries to revert to a prior consistent node $r<t$:
-$\mathrm{BacktrackLocal}(A_{i},\,r):\quad\Phi_{i}^{t}\;\longrightarrow\;\Phi_{i}^{r}.$
-After local rollback, the agent re-evaluates newly arriving evidence.
-Global Conflict and Holistic Updates.
-When contradictions span multiple agents, the Supervisor Agent identifies a minimum conflict set of assertions that must be revised. A global backtracking operation
-$\mathrm{BacktrackGlobal}\bigl{(}\sigma^{t},\,r\bigr{)}$
-reverts all agents from time $t$ to $r$.
+If an internal contradiction $\mathrm{Conflict}_{i}(\Phi_{i}^{t})$ is detected by $A_{V}$, local backtracking is performed to revert $A_{i}$ to a prior consistent node $r<t$:
+$\mathrm{BacktrackLocal}(A_{i},\,r):\Phi_{i}^{t}\longrightarrow\Phi_{i}^{r}.$
+After the local rollback, the agent re-evaluates newly arriving evidence.
+Global Backtracking.
+When contradictions span multiple agents, the $A_{S}$ identifies a minimum conflict set of assertions that must be revised.
+A global backtracking operation: $\mathrm{BacktrackGlobal}\bigl{(}\sigma^{t},\,r\bigr{)}$ reverts all agents from time $t$ to $r$.
 The system discards any statements introduced between $r+1$ and $t$, restoring consistency.
 If the conflict cannot be removed even after a global rollback, the Controller might enforce strategic overrides or disclaim an answer.
-Algorithm 1 summarizes the flow of this multi-agent reversible reasoning design.
-The process starts with question decomposition and sub-question retrieval, followed by verification.
-Local backtracking is invoked as needed to address internal inconsistencies.
-If the conflict persists, it is escalated to the supervisory layer, where partial or holistic backtracking may be applied.
-After conflicts are resolved, the final step integrates all consistent sub-answers.
+The corresponding algorithm is detailed in Appendix E.
+5 Experiments
+5.1 Experimental Setup
+We evaluate ReAgent on three widely used, knowledge-intensive multi-hop QA benchmarks: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), and MuSiQue (Trivedi et al., 2022).
+We compare ReAgent against three main groups of baselines: (1) standard LLMs, (2) dedicated reasoning models, and (3) agent-based models.
+Experimental results demonstrate that ReAgent consistently outperforms all baseline groups and is particularly effective in solving tasks that require iteratively corrected reasoning.
+Datasets.
+HotpotQA is a large-scale, open-domain dataset that explicitly promotes multi-hop reasoning by requiring the integration of information across multiple full-length Wikipedia passages.
+2WikiMultiHopQA selects distinct Wikipedia domains for each question, focusing on evaluating the model’s multi-hop inference over different resources rather than one document.
+Musique is a challenging dataset requiring model’s ability to reason over multiple dispread sentences.
+Following previous work (Trivedi et al., 2023; Press et al., 2023; Gutiérrez et al., 2024), we utilize 1000 random samples from each validation set and corresponding texts as the knowledge base.
+Baselines.
+We meticulously categorize diverse models into different groups to compare their capabilities in multi-hop QA.
+(1) Regular Models: We select non-reasoning models in this group.
+The models include: Llama-4  [^2] https://ai.meta.com/blog/llama-4-multimodal-intelligence/, Qwen-2.5 (Yang et al., 2024a) series, DeepSeek-V3-2024-03 (Liu et al., 2024), Genmini-1.5-Flash-2024-09 (Gemini et al., 2023), Gemini-2.0-flash-2025-02 (Gemini et al., 2023), GPT-4o-latest (Hurst et al., 2024), GPT-4.1 (Meta AI, 2024), GPT-4o with CoT (Wei et al., 2023).
+(2) Reasoning Models: This group includes several strong reasoning baselines for comparison.
+The models include: Qwen-3-Thinking (Qwen3, 2025) in 32B and 253B size, DeepSeek-R1 (DeepSeek-AI, 2025), Gemini-2.5-pro (Google DeepMind, 2025), O1 (OpenAI, 2024), and O3 (OpenAI, 2025).
+(3) Agentic Models.
+Chain-of-Agents (CoA) (Zhang et al., 2024), HippRAG (Gutiérrez et al., 2025), KAG (Liang et al., 2024), and our method.
+To compare fairly, we employ GPT-4o as the backbone for both COT and Agentic Models.
+Implementation.
+For large-scale LLMs (e.g., DeepSeek-V3, the Gemini family, GPT-4o, and Qwen-3-235B), we access the models via API.
+For the medium-sized LLMs, we use their official open-source repositories.
+The temperature is set to $0.3$ to ensure deterministic outputs, while other parameters follow their default settings.
+ReAgent is implemented using GPT-4o as the backbone model.
+Specifically, we set the temperature to $0.8$ for the decomposition agent to encourage diverse reasoning paths, and $0.6$ for all other agents.
+The cost analysis for proprietary models and agentic methods is presented in Appendix B.
+For open-source models, all experiments are conducted using four A100-80G GPUs.
+Metrics. We measure Exact Match (EM) and F1 scores for the QA evaluation.
 Model HotpotQA 2Wiki Musique Average
 EM F1 EM F1 EM F1 EM F1
-Zero-shot
-Qwen-2.5-14b-Instruct 0.205 0.316 0.258 0.308 0.042 0.131 0.168 0.251
-Qwen-2.5-32b-Instruct 0.191 0.304 0.190 0.294 0.041 0.133 0.140 0.243
-Qwen-2.5-72b-Instruct 0.238 0.355 0.286 0.338 0.049 0.141 0.191 0.278
-Llama-3.3-70b-Instruct 0.280 0.415 0.290 0.349 0.063 0.155 0.211 0.306
-DeepSeek-v3 0.293 0.410 0.258 0.332 0.042 0.145 0.197 0.295
-Gemini-1.5-flash 0.236 0.348 0.267 0.311 0.055 0.148 0.186 0.269
-Gemini-2.0-flash 0.317 0.434 0.335 0.391 0.075 0.170 0.242 0.331
-GPT-4o 0.359 0.497 0.346 0.417 0.100 0.217 0.268 0.377
-Knowledge Augmented
-Qwen-2.5-14b-Instruct 0.594 0.768 0.623 0.725 0.305 0.453 0.507 0.648
-Qwen-2.5-32b-Instruct 0.576 0.756 0.537 0.694 0.324 0.460 0.479 0.636
-Qwen-2.5-70b-Instruct 0.604 0.772 0.664 0.753 0.326 0.475 0.531 0.667
-Llama-3.3-70b-Instruct 0.624 0.790 0.698 0.783 0.332 0.470 0.551 0.681
-DeepSeek-v3 0.620 0.792 0.673 0.769 0.368 0.505 0.553 0.688
-Gemini-1.5-flash 0.622 0.784 0.644 0.732 0.314 0.435 0.526 0.680
-Gemini-2.0-flash 0.623 0.790 0.638 0.744 0.306 0.441 0.522 0.685
-GPT-4o 0.612 0.794 0.677 0.780 0.356 0.506 0.548 0.693
+Regular Models
+Llama-4-Scout-17B-16E-Instruct 0.263 0.389 0.332 0.467 0.107 0.185 0.234 0.346
+DeepSeek-V3 0.352 0.491 0.466 0.579 0.223 0.33 0.347 0.467
+Qwen-2.5-32B-Instruct 0.372 0.509 0.557 0.663 0.159 0.273 0.363 0.482
+Qwen-2.5-72B-Instruct 0.363 0.519 0.543 0.631 0.222 0.327 0.376 0.492
+Gemini-1.5-Flash 0.374 0.488 0.563 0.650 0.208 0.310 0.381 0.482
+Gemini-2.0-Flash 0.371 0.490 0.538 0.651 0.246 0.338 0.385 0.493
+GPT-4o 0.381 0.549 0.517 0.649 0.245 0.379 0.381 0.525
+GPT-4.1 0.389 0.563 0.544 0.665 0.271 0.413 0.401 0.547
+CoT (GPT-4o) 0.408 0.531 0.558 0.638 0.272 0.360 0.413 0.509
+Reasoning Models
+Qwen-3-32B-Thinking 0.332 0.474 0.241 0.357 0.387 0.511 0.387 0.511
+DeepSeek-R1 0.356 0.483 0.601 0.707 0.298 0.416 0.418 0.535
+Qwen-3-235B-A22B-Thinking 0.361 0.506 0.624 0.729 0.271 0.387 0.418 0.540
+Gemini-2.5-Pro 0.430 0.560 0.743 0.829 0.383 0.491 0.518 0.626
+O1 0.505 0.661 0.656 0.758 0.417 0.551 0.526 0.656
+O3 0.535 0.696 0.706 0.787 0.442 0.579 0.561 0.687
+Agentic Models (w.GPT-4o)
+CoA Zhang et al. (2024) 0.391 0.558 0.575 0.697 0.239 0.361 0.402 0.539
+HippoRAG (Gutiérrez et al., 2025) 0.528 0.717 0.633 0.725 0.353 0.507 0.504 0.649
+KAG (Liang et al., 2024) 0.603 0.782 0.681 0.781 0.348 0.489 0.544 0.684
 ReAgent (Ours) 0.630 0.795 0.711 0.793 0.371 0.515 0.571 0.701
 Table 1: Performance of different models across three multi-hop QA datasets.
-In each column, the highest performance is highlighted in bold; and within each method group, the top performer is underlined.
-4 Experiments
-4.1 Experimental Setup
-We assessed ReAgent through three widely used knowledge-intensive multi-hop QA datasets: HotpotQA (Yang et al., 2018), 2WikiMultiHotQA (Ho et al., 2020) and Musique (Trivedi et al., 2022).
-We show how ReAgent produces better results compared to baseline models and effectively solves problems requiring continuously corrected reasoning.
-Datasets
-HotpotQA is a large-scaled, open-domain dataset encouraging multi-hop reasoning.
-The dataset provides full Wikipedia passages for information retrieval and integration.
-2WikiMultiHopQA selects distinct Wikipedia domains for each question, focusing on evaluating the model’s multi-hop inference over different resources rather than one document.
-Musique is a challenging dataset requiring model’s ability to reason over multiple dispread sentences.
-Following previous work Trivedi et al. (2023); Press et al. (2023); Gutiérrez et al. (2024), we utilize 1000 random samples from each validation set and corresponding texts as the knowledge base.
-Baselines.
-We include both proprietary and open-sourced LLMs with multiple sizes for evaluation: Qwen-2.5 Yang et al.
-(2024a) of varying parameter scales, the newly Llama-3.3 Meta (2024) and DeepSeek-v3 Liu et al.
-(2024), Gemini Team et al.
-(2023) variants (Gemini-1.5-flash-2024-09, Gemini-2.0-flash-2025-02 as well as GPT-4o-2024-08-06 Hurst et al.
-(2024).
-Settings.
-We hold two distinct settings for the performance evaluation: Zero-shot with CoT only provides the question, expecting to validate LLM’s performance with internal knowledge; Knowledge-Augmented with CoT Liang et al.
-(2024); Gutiérrez et al.
-(2024) setting introduces related documents for reasoning, evaluating model’s capability of information integration especially processing multiple mixed evidence.
-Implementation.
-For super large-size LLMs (DeepSeek-v3, Gemini families, GPT-4o), we access through API, for medium-size LLMs, we access from their official repositories.
-We set the temperature as $0.2$ to get deterministic answers as well as somehow maintain flexibility.
-We utilize the default settings of other parameters.
-ReAgent utilizes GPT-4o as the backbone for execution.
-Specifically, we set the temperature for decomposition agent as $0.8$, the remaining as $0.6$.
-All prompts are available in Appendix A.
-Metrics.
-We measure Exact Match (EM) and F1 scores for the QA evaluation.
+In each column, the highest and the second highest performance is highlighted in red and blue; and within each method group, the top performer is underlined.
 5.2 Results and Analysis
-Table 1 illustrates the comprehensive performance.
-ReAgent outperforms all the baseline models in both metrics across all datasets.
-Specifically, our framework achieves an average EM of 0.571 and an average F1 of 0.701, outperforming GPT-4o by 2.3% in EM and 0.8% in F1.
-External Knowledge Matters.
-Definitely, we observe that zero-shot performance is substantially lower, reflecting the difficulty of multi-hop reasoning without relevant background.
-With intensive knowledge, all models outperform their zero-shot versions in different levels, presenting their internal reasoning abilities.
-Scaling-law not always works.
-We carefully select baseline models across small(14b), medium(70b), large(>100b) size for comparisons.
-Interestingly, the results show that smaller models sometimes outperform larger models.
-For example, Qwen-2.5-14b model generally defeats its 32b counterparts in both zero-shot and knowledge augmented settings.
-Similarly, Llama-3.3 approach or even surpass the super large-scale models (e.g.
-Gemini-1.5-flash and DeepSeek-v3 in zero-shot, Gemini-2.0-flash and GPT-4o in knowledge-intensive setting).
-This suggests that multi-hop QA can be aided by carefully tuned architectures, especially if they incorporate chain-of-thought or knowledge-aware strategies.
+Table 1 presents the overall performance across all datasets.
+ReAgent consistently outperforms all baseline models on both EM and F1 metrics.
+Specifically, it achieves an average EM of 0.571 and an average F1 of 0.701, surpassing the strongest baseline—knowledge-augmented GPT-4o—by 2.3% in EM and 0.8% in F1.
+It also outperforms strong reasoning models, including O1 and O3.
+Moreover, it also surpasses the recent agentic models such as CoA, HippoRAG, and KAG across all datasets.
+We summarize the following insights:
+Agent-based and reasoning enhances performance.
+Some reasoning models and agentic models have a competitive performance (such as O1, O3, and KAG) than regular models.
+Notably, ReAgent outperforms O3, one of the strongest reasoning models, and improves significantly over GPT-4o+CoT in both EM and F1.
+This suggests that modular execution and collaboration among agents provide a clearer advantage in complex, multi-hop settings.
+While ReAgent can theoretically use stronger backbones like O1 or O3, their high cost (645\$ and 546\$, respectively, as shown in Appendix B) makes them not a good option.
+Our focus is to show that even with GPT-4o, ReAgent is able to beat the expensive O1 and O3.
+ReAgent excels in manageable context length.
+Specifically, ReAgent achieves the highest F1 score on HotpotQA (0.795), demonstrating strong performance in tasks that require structured reasoning over relatively short contexts.
+In contrast, Gemini-2.5-Pro performs well on 2Wiki but underperforms on other datasets, while O1 and O3 show stronger results on Musique but fall short on HotpotQA.
+This discrepancy is partly attributed to the increased complexity of the Musique dataset, which involves longer contexts.
+While ReAgent is built on GPT-4o and may be less optimized for long documents than dedicated reasoning models, its agentic design, including traceback and self-check mechanisms, proves particularly effective in scenarios requiring precision, stepwise planning, and robust verification.
+These strengths enable ReAgent to outperform other models on tasks like HotpotQA, highlighting its superior general reasoning capabilities within manageable context lengths.
 Overall, we emphasized the value of reversible reasoning mechanisms to mitigate error propagation.
-The data in Table 1 align with this assumption: even with strong models, single-pass reasoning can fail unless the correct context is identified or re-checked.
+Results in Table 1 align with this assumption: even with strong models, single-pass reasoning can fail unless the correct context is identified or re-checked.
 Our method’s ability to backtrack helps resolve contradictions, leading to more stable performance in multi-hop scenarios.
-Figure 4: Ablation Study Results Comparison.
 5.3 Ablation Study
-To gauge the importance of the backtracking mechanism, we conduct an ablation study by disabling it under the same settings.
-Specifically, the system then proceeds in a strictly forward fashion, without the ability to revert to earlier intermediate states.
-Fig.4 compares performance between different backbones, with and without backtracking, on three datasets.
-As shown, the performance drop highlights the importance of backtracking in mitigating error propagation, where an early misstep without backtracking cannot be corrected, resulting in deteriorated performance.
+Effectiveness Analysis of Backtracking Mechanism.
+We conduct an ablation study by disabling backtracking under the same settings to verify its importance.
+In this setup, the system operates strictly forward without the ability to revert to earlier intermediate states.
+Figure 3 compares performance between different backbones, with and without backtracking, on HotpotQA dataset.
+The performance drop highlights the importance of backtracking in mitigating error propagation, where an early misstep without backtracking cannot be corrected, resulting in deteriorated performance.
 Notably, the DeepSeek-V3 backbone exhibits general improvements with backtracking, demonstrating its robustness across various settings.
+Impact of Different Local Backtracking (BT) Depth.
+To further analyze the impact of local backtracking depths (the number of steps the system is allowed to revert), we analyze the performance of our proposed ReAgent under different settings, as shown in Figure 4 (left).
+The results show that on two selected backbones, DeepSeek-V3 and GPT-4o, the performance of ReAgent improves as the local backtracking depth increases, suggesting that a deeper backtracking depth benefits the agent by enabling it to effectively recover from earlier reasoning errors.
+However, the benefits gradually saturate, indicating that further increasing the backtracking depth yields diminishing returns.
+Figure 3: Ablation Study on Local and Global Backtracking (BT): EM comparison on HotpotQA using GPT-4o and DeepSeek-V3.
+Impact of the Number of Decomposed Sub-Questions.
+Figure 4 (right) demonstrates how a different number of decomposed sub-questions affects performance.
+The results show that decomposing the input question into multiple sub-questions significantly improves performance.
+Specifically, using 3 sub-questions is the optimal trade-off choice, with an improvement of 0.161 EM on the GPT-4o and 0.142 EM on DeepSeek-V3 compared to directly processing a single question.
+Beyond this point, increasing the number of sub-questions brings only marginal gains but higher costs.
+Figure 4: Ablation Study on Backtracking Depth (left) and Number of Decomposed Sub-Questions (right): EM comparison on HotpotQA using GPT-4o and DeepSeek-V3.
+Figure 5: Case study comparing GPT-O3 (left), CoA (middle), and ReAgent (right) on HotpotQA.
+The back-tracking mechanism enhances iterative reasoning, enabling conflict detection and correction to reach the correct answer.
 6 Case Study
-We illustrate our multi-agent backtracking (ReAgent) through a single-culprit puzzle with four suspects {A, B, C, D}.
-Each suspect gives statements about who might be guilty or lying.
-The puzzle’s only correct solution is that B is the culprit, but identifying this requires partially retracting initial assumptions along the way.
-Detailed puzzle statement can be found in Appendix B.
-Why a Rollback Mechanism is Needed.
-Naive single-pass (or single-thread) models often fixate prematurely on one suspect, disregard contradictory evidence, and produce unsound conclusions.
-By contrast, our multi-agent approach enumerates possible culprits in parallel, detects conflicts, and rolls back to revise incorrect assumptions.
-Below is a simplified agentic workflow capturing the time steps $T_{1}$ through $T_{7}$ during which each agent processes statements, detects a conflict, and reverts.
-Algorithm 2 ReAgent Multi-Agent Puzzle Solving
-1:Input: Puzzle statements from A, B, C, D.
-2:Goal: Identify unique culprit (one of {A,B,C,D}) with minimal contradictions.
-3:Initialization
-[…]
-[…]
-6:Round 1: Checking A as culprit
-7:$T_{1}:$ $\text{Checker}_{A}$:  Evaluate puzzle statements assuming $A$ is culprit.
-8:$T_{2}:$ $\text{Checker}_{A}$ detects contradiction:  ($A$’s claims vs. $C$’s claims cannot both hold).
-9:$T_{3}:$ Conflict Detector signals rollback to discard $A$-culprit assumption.
-10:Round 2: Checking B as culprit
-11:$T_{4}:$ $\text{Checker}_{B}$ collects statements $\{A,B,C,D\}$ under $B$=culprit.
-12:$T_{5}:$ Finds no fatal conflicts (B’s statements can be partly false, others partly true).
-13:$T_{6}:$ Conflict Detector sees consistency, no rollback needed.
-14:Rounds 3 and 4: Checking C or D as culprit
-15:$T_{7}:$ Similar to $H_{A}$, each leads to irreconcilable contradictions.
-16:          $\Rightarrow$ Supervisor triggers rollback again; discards these.
-17:Conclusion:  Only $H_{B}$ remains consistent throughout.  Answer: $B$ is culprit.
-Discussion.
-This staged procedure allows each culprit hypothesis to be explored (lines 7–13 of Algorithm 2).
-Any contradictions are escalated at $T_{2}$ or $T_{7}$ to the Conflict Detector, which triggers a rollback to the prior consistent node.
-Ultimately, only the assignment “$B$ is culprit” withstands these iterative checks.
-Without rollback, a linear solver often locks in a wrong suspect early and fails to correct itself.
-For extended discussions of single-thread LLM failures, and the complete multi-agent transcripts, see Appendix B and  C.
-6 Related Work
-6.1 Prompting and Iterative Reasoning
-Large Language Models (LLMs) can solve complex tasks by leveraging prompting strategies that encourage step-by-step reasoning, often referred to as chain-of-thought (CoT) prompting Wei et al.
-(2022), which has demonstrated improved performance in arithmetic, commonsense, and symbolic reasoning tasks, particularly in sufficiently large models.
-Sampling multiple reasoning paths and selecting the most consistent answer, known as self-consistency, further boosts accuracy by treating final answers as a “majority vote” across different solutions Wang et al.
-(2023).
-To reduce the need for hand-tuned prompts, automated prompt engineering techniques have been explored.
-One approach, Automatic Prompt Engineer (APE), treats prompt generation as a search problem: one LLM proposes candidate instructions while another evaluates them, allowing the system to refine prompts over multiple tasks Zhou et al.
-(2023).
-LLMs can thus act as both reasoning engines and prompt designers.
-Beyond one-shot responses, iterative refinement frameworks empower LLMs to correct and enhance their outputs.
-For instance, the ReAct framework interleaves reasoning traces with actions (e.g., external API calls), enabling LLMs to incorporate feedback and reduce factual errors Yao et al.
-(2023).
-Other methods like Reflexion Shinn et al.
-(2023) and Self-Refine Madaan et al.
-(2023) introduce self-feedback loops, where a model critiques its own response and revises it in subsequent attempts, leading to more reliable and transparent solutions.
-Prompting methods have also expanded into multimodal settings, where models process both text and visual inputs.
-In Multimodal-CoT, the system first generates a textual chain-of-thought integrating visual information, then produces an answer from that combined rationale Zhang et al.
-(2023).
-6.2 Multi-Agent Collaboration, Debate, and Scalability
-Recent work explores multi-agent approaches, where multiple LLMs (or instances of the same model) collaborate or compete toward a shared goal.
-For instance, the CAMEL framework assigns different roles (e.g., “user” and “assistant”) to LLM agents that communicate in a dialogue-based fashion, breaking down tasks more systematically Li et al.
-(2023).
-Role specialization allows agents to focus on different aspects of the problem, and voting or consensus mechanisms—akin to self-consistency—can yield more robust answers.
-Debate-oriented methods push multi-agent interactions further by having agents argue for different viewpoints.
-After hearing opposing arguments, a human or model judge selects the most convincing solution Khan et al.
-(2024).
-Similarly, in an interconsistency framework, multiple agents provide arguments and counterarguments over multiple rounds, eventually reaching a consensus Xiong et al.
-(2023).
-Such adversarial or critical exchanges can improve truthfulness and correctness, though an overly dominant model can bias outcomes unless moderated by a sufficiently capable judge Xiong et al.
-(2023).
-These multi-agent and multimodal systems introduce scalability and coordination challenges.
-As more agents or steps are added, communication overhead can grow significantly, making real-time collaboration difficult.
-Research highlights the need for efficient structures—e.g., hierarchical controllers or learned protocols—to balance the benefits of specialized agents with the system’s complexity Li et al.
-(2023); Xiong et al.
-(2023).
+6.1 ReAgent Walk Through
+To illustrate how recursive feedback enables robust multi-agent reasoning, we present a case where the system answers the question: “Which U.S.
+state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?”, shown in Figure 6.
+The question is first decomposed by the Question Decomposer agent ($A_{Q}$) into four sub-questions: identifying the host state of the 1984 Olympics, retrieving its capital and largest city, comparing their populations, and returning the state if the capital is smaller.
+The Retriever agent ($A_{R}$) retrieves that California hosted the 1984 Olympics, with Sacramento as its capital and Los Angeles as its largest city.
+However, inconsistent population data for Sacramento (508k vs.
+1.5M) triggers a local conflict, which is resolved by the Verifier agent ($A_{V}$) through local backtracking—discarding the unreliable 1.5M estimate.
+In this case, the local backtracking step is able to figure out the correct information, then the Answer Assembler ($A_{A}$) is able to gather all the information locally and then globally, and finally present the final answer to be California.
+We eliminate other parts when another conflict may occur, a full walk-through is presented in Appendix D.
+Another case study is presented in Appendix C
+6.2 Comparison with Baseline Models
+To elucidate differences in reasoning dynamics, we perform a case study comparing ReAgent with a single-agent reasoning baseline, O3 and an agentic baseline, Chain-of-Agents (CoA) Zhang et al.
+(2024) on the query ‘‘What is the name of the fight song of the university whose main campus is in Lawrence, Kansas and whose branch campuses are in the Kansas City metropolitan area?’’, as illustrated in Figure 7.
+The full comparison is presented in Appendix D.
+The example question requires two separate hops: (i) identifying which university satisfies the geographical constraints, and (ii) retrieving the “official” fight-song of that school.
+Although seemingly simple, the task hides a subtle distinction: the University of Kansas (KU) has both an official fight song (“Kansas Song”, adopted in 1928) and a far better-known rally tune (“I’m a Jayhawk”).
+Correctly answering therefore hinges on (a) verifying that KU is the only university matching the campus pattern, and (b) resolving the potential conflict between the two candidate songs.
+The reasoning model O3 (left) performs the first hop successfully but then falls back on memorised popularity cues, assuming that the song most familiar to the public must be official.
+The CoA (middle) multi-worker system executes three independent workers and lets a manager pick the majority vote.
+Worker 1 fails to locate any fight-song information due to the chunk division.
+Worker 2 repeats O3’s mistake.
+Worker 3 extracts the correct answer “Kansas Song” but discards it because it is less commonly used, and the third fails to locate any fight-song information.
+The manager simply adopts the majority answer (“I’m a Jayhawk”), which causes the error.
+Our ReAgent (right) framework handles the same example with six specialized retriever agents answering the sub-questions generated by the Decomposer: (a) Q1–Q3 (entity verification).
+Three agents collectively confirm that only KU satisfies the main-campus/branch-campus constraints, eliminating alternative schools early.
+(b) Q4–Q5 (conflict surfacing).
+Independent retrieval agents surface both “I’m a Jayhawk” and “Kansas Song”, triggering an explicit conflict state.
+(c) Backtracking (Q6).
+The controller reverses to Q4 and re-weights evidence.
+The source for “Kansas Song” contains the enactment year (1928) and the keyword “official”, while “I’m a Jayhawk” is marked only as a “rally tune”.
+Therefore, our model successfully reach the correct answer “Kansas Song”.
+This demonstrates that, thanks to its conflict-detection and backtracking mechanisms, ReAgent can revisit earlier reasoning, resolve contradictory evidence, and consistently converge on the correct knowledge-grounded answer.
+Figure 6: Case study: We illustrate the main steps of how ReAgent answers a question, where local backtracking is highlighted to resolve a conflict.
 7 Conclusion
@@
 Limitations
-Despite its advantages, our framework inherits several constraints:
-  • Complex Coordination: Multi-agent interaction and backtracking mechanisms introduce additional computational overhead and higher implementation complexity.
-  • Dependence on Retrieval Quality: As with many QA systems, our accuracy hinges on retrieving relevant and sufficiently detailed evidence.
-Misleading or scarce information can impair performance, even with backtracking.
-  • Scalability to Larger Contexts: While our approach manages conflicts through localized rollback, handling very large or highly diverse corpora may require more sophisticated retrieval and aggregation strategies to maintain efficiency.
-  • Model Assumptions: We rely on the underlying language models’ capacity for reliable verification.
-If the base model’s reasoning is systematically biased, backtracking alone may not fully correct errors.
-Ethic Statement
-Our method focuses on improving the interpretability and robustness of multi-hop question answering rather than amplifying any content that could be personally or socially harmful.
-We do not incorporate user-sensitive or private data, and the knowledge sources employed are publicly accessible.
-However, as with any automated system, inaccurate or biased inputs may lead to flawed inferences if not properly filtered or monitored.
-We encourage careful data curation and ongoing human oversight to mitigate unintended consequences.
-Furthermore, we recommend that users verify critical or sensitive outputs, since backtracking techniques reduce—but cannot entirely eliminate—residual errors inherent in large language models.
+In this paper, we present the ReAgent model for multi-hop QA, which introduces explicit backtracking mechanisms that allow the system to correct errors during reasoning, leading to more accurate and reliable answers.
+However, this design introduces certain limitations.
+First, the ability to backtrack multiple times can increase the overall inference time, potentially making the reasoning process less efficient.
+Second, although ReAgent demonstrates improved performance over selected reasoning models, its more complex architecture may reduce its robustness in scenarios with limited resources or noisy inputs.
+In the future, we aim to improve the design of the agents, with a particular focus on enhancing their collaboration strategies to further reduce error propagation and improve reasoning efficiency.
 Appendix
@@
 Downstream agents read the previous agent’s JSON fields directly, enabling a structured, chain-of-thought style pipeline that remains reversible at every step.
-Appendix B Case Study
-B.1 Puzzle Statement
+Appendix B Analysis of Computational Cost
+We conduct a comparative analysis of computational cost on the HotpotQA, 2Wiki, and MuSiQue datasets across GPT-4o, reasoning models O1 and O3, and our ReAgent.
+The evaluation includes average inference calls, inference time (s), input/output tokens, and total cost(\$), as shown in Table 2.
+The results highlight that our ReAgent, built on GPT-4o, achieves superior performance compared to reasoning models O1 and O3, while incurring significantly lower computational cost.
+Although its cost is higher than the GPT-4o baseline, ReAgent delivers substantially better performance.
+These findings demonstrate the effectiveness of our method in balancing performance and efficiency.
+Model Avg. Avg. Avg. Avg. Total
+Calls Time(s) Input(T) Output(T) Cost(\$)
+GPT-4o 1 2.7 9 289 5  69
+O1 1 43 9 289 1 471  645
+O3 1 27 9 289 1 471  546
+Ours (ReAgent,(GPT-4o)) 29 46 12 400 1 920 275
+Table 2: Cost Comparison.
+T denotes Tokens.
+Note that the input token counts are aggregated across the three datasets.
+Since the complete reasoning outputs for O1 and O3 were unavailable, their output token counts are estimated based on the reasoning outputs generated by Qwen-3-235B.
+Appendix C Case Study on Puzzle Solving
+Algorithm 1 ReAgent Multi-Agent Puzzle Solving
+1:Input: Puzzle statements from A, B, C, D.
+2:Goal: Identify unique culprit (one of {A,B,C,D}) with minimal contradictions.
+3:Initialization
+[…]
+[…]
+6:Round 1: Checking A as culprit
+7:$T_{1}:$ $\text{Checker}_{A}$:  Evaluate puzzle statements assuming $A$ is culprit.
+8:$T_{2}:$ $\text{Checker}_{A}$ detects contradiction:  ($A$’s claims vs. $C$’s claims cannot both hold).
+9:$T_{3}:$ Conflict Detector signals rollback to discard $A$-culprit assumption.
+10:Round 2: Checking B as culprit
+11:$T_{4}:$ $\text{Checker}_{B}$ collects statements $\{A,B,C,D\}$ under $B$=culprit.
+12:$T_{5}:$ Finds no fatal conflicts (B’s statements can be partly false, others partly true).
+13:$T_{6}:$ Conflict Detector sees consistency, no rollback needed.
+14:Rounds 3 and 4: Checking C or D as culprit
+15:$T_{7}:$ Similar to $H_{A}$, each leads to irreconcilable contradictions.
+16:          $\Rightarrow$ Supervisor triggers rollback again; discards these.
+17:Conclusion:  Only $H_{B}$ remains consistent throughout.  Answer: $B$ is culprit.
+We illustrate our multi-agent backtracking (ReAgent) through a single-culprit puzzle with four suspects {A, B, C, D}.
+Each suspect gives statements about who might be guilty or lying.
+The puzzle’s only correct solution is that B is the culprit, but identifying this requires partially retracting initial assumptions along the way, as shown in Algorithm 1.
+Why a Rollback Mechanism is Needed.
+Naive single-pass (or single-thread) models often fixate prematurely on one suspect, disregard contradictory evidence, and produce unsound conclusions.
+By contrast, our multi-agent approach enumerates possible culprits in parallel, detects conflicts, and rolls back to revise incorrect assumptions.
 Suspects and Rules:
@@
 Which single suspect is guilty, respecting the puzzle rules?
-B.2 Agents and Their Roles
+Appendix D Case Study on Traditional Multi-hop QA Tasks
+D.1 User’s question:
+"Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?"
+To answer this, the system must:
+  1. 1.
+    Identify the state that hosted the 1984 Summer Olympics.
+  2. 2.
+    Compare the capital city’s population to the largest city’s population in that state.
+  3. 3.
+    Confirm the capital city’s population is indeed smaller.
+  4. 4.
+    Provide the name of that state.
+Although it may appear straightforward, we deliberately introduce contradictory or erroneous data along the way to showcase the reversible (backtracking) mechanisms.
+D.2 Agents and Their Roles
   1. 1.
@@
     Provides strategic oversight. In the example below, it will issue “challenges” to specific assertions when local backtracking fails.
-B.3 Walkthrough of the Reasoning and Backtracking
-B.3.1 Initial Question Decomposition
+D.3 Walkthrough of the Reasoning and Backtracking
+Initial Question Decomposition
   1. 1.
@@
     These sub-questions ($q_{1}$$q_{4}$) are broadcast to the Retriever ($A_{R}$) and Verifier ($A_{V}$).
-B.3.2 Retrieval and a Local Conflict
+Retrieval and a Local Conflict
   1. 1.
@@
 (This demonstrates how a single agent can retract contradictory or low-confidence evidence without halting the entire process.)
-B.3.3 Assembling Partial Answers & Uncovering a Global Conflict
+Assembling Partial Answers & Uncovering a Global Conflict
   1. 1.
@@
     These two are clearly incompatible. The next step is to identify which statement should be retracted system-wide.
-B.3.4 Global Backtracking and Final Resolution
+Global Backtracking and Final Resolution
   1. 1.
@@
     Final Answer: California.
-B.4 Key Observations
+D.4 Full Comparison with Baseline Models (Figure 7)
+Figure 7: Full Comparison with Baseline Models
+D.5 Key Observations
   • Early-Stage Conflict Resolution: The Verifier Agent ($A_{V}$) performed local backtracking when it discovered contradictory population data for Sacramento.
@@
   • Strategic Re-check and Override: The Controller Agent ($A_{C}$) could forcibly challenge the suspicious statement “Capital(California, Los Angeles)” to eliminate it from the knowledge pool, thus preserving the correct solution.
+Algorithm 2 Multi-Agent Reversible Reasoning
+1:$Q_{0}$: Main question
+$A_{Q}.\mathrm{decompose}(Q_{0})\to\{q_{1},q_{2},\dots\}$
+3:Broadcast $\mathrm{assert}(q_{i})$ to $A_{R},A_{V},A_{A}$
+4:for each $q_{i}$ in $\{q_{1},q_{2},\dots\}$ concurrently do
+5:     $E\leftarrow A_{R}.\mathrm{retrieve}(q_{i})$
+6:     $A_{V}.\mathrm{verify}(E)$
+7:     if $A_{V}$ detects conflict locally then
+8:         $A_{V}.\mathrm{BacktrackLocal}(r_{V})$
+9:         if conflict persists then
+10:              raise Conflict to $A_{S}$
+11:         end if
+12:     end if
+13:     $A_{A}.\mathrm{storePartialAnswer}(q_{i},E)$
+14:end for
+15:if Global conflict signaled then
+16:     $A_{S}.\mathrm{HolisticUpdate}(\Pi_{j})$
+17:     if $A_{C}$ intervention needed then
+18:         $A_{C}.\mathrm{challenge}(\varphi)$ or override
+19:     end if
+20:end if
+21:Final $\leftarrow A_{A}.\mathrm{assembleAnswer}(\{q_{1},q_{2},\dots\})$
+22:return Final Answer
 Overall, this case exemplifies how reversible, multi-hop reasoning allows for robust error correction: local invalid data is undone swiftly, while deeper logic conflicts trigger system-wide backtracking.
 Consequently, the answer “California” remains stable and correct, ensuring that a single faulty retrieval step does not irreversibly corrupt the entire reasoning chain.
-Appendix C Traditional Multi-hop QA Tasks
-1. The Question
-C.0.1 User’s question:
-"Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?"
-To answer this, the system must:
-  1. 1.
-    Identify the state that hosted the 1984 Summer Olympics.
-  2. 2.
-    Compare the capital city’s population to the largest city’s population in that state.
-  3. 3.
-    Confirm the capital city’s population is indeed smaller.
-  4. 4.
-    Provide the name of that state.
-Although it may appear straightforward, we deliberately introduce contradictory or erroneous data along the way to showcase the reversible (backtracking) mechanisms.
-C.0.2 Agents and Their Roles
-  1. 1.
-    Question-Decomposer Agent ($A_{Q}$)
-    Splits the complex question into sub-questions:
-    • Sub-question 1: Which U.S. state hosted the 1984 Summer Olympics?
-    • Sub-question 2: What is the capital city of that state, and what is the largest city?
-    • Sub-question 3: Compare the populations of the capital and the largest city.
-    • Sub-question 4: Return the state name if the capital’s population is smaller.
-  2. 2.
-    Retriever Agent ($A_{R}$)
-    Searches external knowledge (e.g., a text corpus or knowledge graph) to gather relevant facts:
-    • Text passages or data about U.S. states, capitals, largest cities, and historical Olympic hosts.
-  3. 3.
-    Verifier Agent ($A_{V}$)
-    Cross-checks newly retrieved information for consistency against its local knowledge or prior verified facts. If a contradiction is detected, it triggers local backtracking.
-  4. 4.
-    Answer-Assembler Agent ($A_{A}$)
-    Synthesizes partial answers from the other agents. If contradictory partial answers cannot be reconciled at the local level, $A_{A}$ escalates the conflict to the Supervisory layer.
-  5. 5.
-    Supervisor Agent ($A_{S}$)
-    Oversees system-wide conflicts. It can perform global backtracking (rolling back all agents) if needed.
-  6. 6.
-    Controller Agent ($A_{C}$)
-    Provides strategic oversight. In the example below, it will issue “challenges” to specific assertions when local backtracking fails.
-C.0.3 Walkthrough of the Reasoning and Backtracking
-3.1 Initial Question Decomposition
-  1. 1.
-    $A_{Q}$ receives the question:
-    $A_{Q}$ parses the main query:
-    “Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?”
-  2. 2.
-    $A_{Q}$ produces sub-questions:
-    • $q_{1}$: “Which U.S. state hosted the 1984 Summer Olympics?”
-    • $q_{2}$: “What is the capital city of that state? Also, what is the largest city of that state?”
-    • $q_{3}$: “Compare the population of the capital city to that of the largest city.”
-    • $q_{4}$: “Return the state if the capital’s population is smaller.”
-    These sub-questions ($q_{1}$$q_{4}$) are broadcast to the Retriever ($A_{R}$) and Verifier ($A_{V}$).
-3.2 Retrieval and a Local Conflict
-  1. 1.
-    $A_{R}$ retrieves answers:
-    • For $q_{1}$, $A_{R}$ returns:
-      “The 1984 Summer Olympics were primarily hosted in Los Angeles, California.”
-    • For $q_{2}$, $A_{R}$ retrieves partial data:
-      • Capital of California is Sacramento.
-      • Largest city is Los Angeles.
-      • Population estimates (unfortunately, one item is erroneous or inconsistent):
-          Sacramento population: $\sim 508,000$ (from official records).
-          Los Angeles population: $\sim 3,900,000$ (correct).
-          But a second retrieved record incorrectly lists Sacramento’s population as 1,500,000.
-  2. 2.
-    $A_{V}$ (Verifier) checks local consistency:
-    $A_{V}$ notices that two different population figures for Sacramento are contradictory:
-    • Data point $D_{1}$: 508,000
-    • Data point $D_{2}$: 1,500,000
-    $A_{V}$ identifies these as mutually exclusive facts. It flags a local conflict:
-    “Population of Sacramento cannot be both 508k and 1.5M.”
-  3. 3.
-    Local Backtracking by $A_{V}$:
-    Before finalizing any partial answer, $A_{V}$ backtracks to the checkpoint prior to adopting the second, suspicious data point $D_{2}$.
-    $A_{V}$ discards the contradictory population figure (1.5M) and retains only 508k as the consistent number for Sacramento.
-    The system continues forward with the corrected value for Sacramento’s population: $\sim 508,000$.
-(This demonstrates how a single agent can retract contradictory or low-confidence evidence without halting the entire process.)
-3.3 Assembling Partial Answers & Uncovering a Global Conflict
-  1. 1.
-    $A_{A}$ (Answer-Assembler) integrates partial conclusions:
-    So far, the state that hosted the 1984 Olympics is California.
-    The capital is Sacramento (population: $\sim 508,000$).
-    The largest city is Los Angeles (population: $\sim 3,900,000$).
-  2. 2.
-    $A_{A}$ composes a preliminary final:
-    Since Sacramento’s population ($\sim 508,000$) is indeed smaller than Los Angeles’s ($\sim 3,900,000$), the preliminary answer is: “California” should be the correct state.
-  3. 3.
-    New conflict introduced:
-    However, suppose the Retriever ($A_{R}$)—in parallel—fetched an alternative “capital city” record stating that Los Angeles was once referred to as the “capital” in a historical context (erroneous snippet from a non-authoritative source).
-    This implies the contradictory statement: “Los Angeles is also the capital of California,” which directly conflicts with the known fact “Sacramento is the capital of California.”
-  4. 4.
-    $A_{V}$ tries local resolution:
-    $A_{V}$ cannot reconcile “Los Angeles is capital” with “Sacramento is capital.”
-    Because each piece of evidence was introduced into separate sub-threads, a single local reversion inside $A_{V}$ might not suffice.
-The conflict is also recognized by $A_{A}$ when finalizing sub-answers.
-  5. 5.
-    Conflict escalates to the Supervisor ($A_{S}$):
-    As soon as multiple agents disagree over fundamental facts (the identity of the capital), the system triggers a global conflict signal.
-    $A_{S}$ identifies the minimal conflicting set:
-    • Capital(California, Sacramento)
-    • Capital(California, Los Angeles)
-    These two are clearly incompatible. The next step is to identify which statement should be retracted system-wide.
-3.4 Global Backtracking and Final Resolution
-  1. 1.
-    Global Backtracking:
-    The Supervisor Agent ($A_{S}$) issues a global backtracking command, rolling the entire system’s knowledge to a shared checkpoint before the contradictory capital reference was accepted.
-    All local knowledge caches revert to a consistent state in which “Sacramento is the capital” is still true, and “Los Angeles is the capital” is no longer present.
-  2. 2.
-    Controller ($A_{C}$) challenges the suspicious assertion:
-    To prevent the same contradiction from reappearing, $A_{C}$ explicitly “challenges” the statement Capital(California, Los Angeles).
-    This statement is reevaluated or ignored based on domain knowledge or reliability checks (e.g., confidence weighting from the retriever and the verifier).
-    The system confirms that Los Angeles was never the official capital.
-  3. 3.
-    Assemble the final consistent answer:
-    With the conflicting statement removed, the pipeline reaffirms:
-    1. (i)
-      The 1984 Summer Olympics took place in Los Angeles, California.
-    2. (ii)
-      California’s capital is Sacramento.
-    3. (iii)
-      Sacramento’s population ($\sim 508,000$) is smaller than Los Angeles’s ($\sim 3,900,000$).
-    Final Answer: California.
-C.0.4 4. Key Observations
-  • Early-Stage Conflict Resolution: The Verifier Agent ($A_{V}$) performed local backtracking when it discovered contradictory population data for Sacramento.
-This promptly removed an incorrect data point without halting the entire process.
-  • Escalation of Irreconcilable Contradictions: When two different sub-threads provided fundamentally clashing information (competing claims about the capital), the system automatically escalated the conflict to the Supervisor ($A_{S}$) for global action.
-  • Strategic Re-check and Override: The Controller Agent ($A_{C}$) could forcibly challenge the suspicious statement “Capital(California, Los Angeles)” to eliminate it from the knowledge pool, thus preserving the correct solution.
-Overall, this case exemplifies how reversible, multi-hop reasoning allows for robust error correction: local invalid data is undone swiftly, while deeper logic conflicts trigger system-wide backtracking.
-Consequently, the answer “California” remains stable and correct, ensuring that a single faulty retrieval step does not irreversibly corrupt the entire reasoning chain.
+Appendix E Reversible Reasoning Algorithm
+Algorithm 2 summarizes the flow of this multi-agent reversible reasoning design.
+The process starts with question decomposition and sub-question retrieval, followed by verification.
+Local backtracking is invoked as needed to address internal inconsistencies.
+If the conflict persists, it is escalated to the supervisory layer, where partial or holistic backtracking may be applied.
+After conflicts are resolved, the final step integrates all consistent sub-answers.